### v1

# GaussianOcc: Fully Self-supervised and Efficient 3D Occupancy Estimation with Gaussian Splatting

Abstract
We introduce GaussianOcc, a systematic method that investigates the two usages of Gaussian splatting for fully self-supervised and efficient 3D occupancy estimation in surround views. First, traditional methods for self-supervised 3D occupancy estimation still require ground truth 6D poses from sensors during training. To address this limitation, we propose Gaussian Splatting for Projection (GSP) module to provide accurate scale information for fully self-supervised training from adjacent view projection. Additionally, existing methods rely on volume rendering for final 3D voxel representation learning using 2D signals (depth maps, semantic maps), which is both time-consuming and less effective. We propose Gaussian Splatting from Voxel space (GSV) to leverage the fast rendering properties of Gaussian splatting. As a result, the proposed GaussianOcc method enables fully self-supervised (no ground truth pose) 3D occupancy estimation in competitive performance with low computational cost (2.7 times faster in training and 5 times faster in rendering). The relevant code will be available in https://github.com/GANWANSHUI/GaussianOcc.git.
Figure 1: Problem setting of GaussianOcc. Given the surround image sequence and its corresponding 2D semantic annotation, GaussianOcc is able to perform 3D occupancy estimation without the need for the ground truth occupancy label and ground truth 6D pose for training.
1 Introduction
In recent years, both vision-centric and grid-centric perception approaches have gained significant attention in the autonomous driving industry and academia [35, 39]. Among these, surround view 3D occupancy estimation [32, 45, 51, 48] has emerged as a core perception task and a promising alternative to bird’s-eye view (BEV) methods [9, 29, 30]. To facilitate 3D occupancy estimation, several benchmarks have been developed for supervised training [45, 43, 41, 42], though these require substantial effort in 3D annotation. To reduce the burden of annotation, self-supervised [12, 54, 24, 5, 18] and weakly-supervised [37] learning approaches based on volume rendering have been proposed [36, 46, 11]. Volume rendering allows 3D representation learning using 2D supervision signals, such as 2D semantic maps and depth maps, thereby eliminating the need for extensive 3D annotation during training.
The existing methods [54, 24] achieve self-supervised learning through volume rendering, where the 2D semantic map supervision is derived from open-vocabulary semantic segmentation [55], and the depth map supervision is obtained from self-supervised depth estimation [13]. However, these approaches face two significant limitations. First, volume rendering is performed at real-world scale, which requires the availability of ground truth 6D pose to calculate the multi-view photometric loss across sequential images. Second, the inefficiency in volume rendering poses a challenge, the same as in novel view synthesis tasks [27, 1, 19], due to the dense sampling operation required. These limitations impede the development of a more general and efficient paradigm for self-supervised 3D occupancy estimation.
To address the aforementioned limitations, we explore a fully self-supervised and efficient approach to 3D occupancy estimation using Gaussian splatting [27, 1]. Specifically, we introduce the use of Gaussian splatting to perform cross-view splatting, where the rendered image constructs a cross-view loss that provides scale information during joint training with the 6D pose network. This eliminates the need for a ground truth 6D pose during training. To improve rendering efficiency, we move away from the dense sampling required in traditional volume rendering. Instead, we propose performing Gaussian splatting directly from the 3D voxel space. In this approach, each vertex in the voxel grid is treated as a 3D Gaussian, and we directly optimize the attributes of these Gaussians—such as semantic features and opacity—within the voxel space. Through this novel approach, our proposed method, GaussianOcc, makes a progress toward fully self-supervised and efficient 3D occupancy estimation, as outlined in Figure 1.
The main contributions of this work are summarized as follows.
  • We introduce the first fully self-supervised method for efficient surrounding-view 3D occupancy estimation, featuring the exploration of the Gaussian splatting.
  • We propose Gaussian splatting for cross-view projection, which could provide the scale information to get rid of the need of a ground truth 6D pose during the training.
  • We propose Gaussian splatting from voxel space achieving competitive performance with 2.7 times faster training and 5 times faster rendering compared with the volume rendering.
2 Related work
2.1 Surround view depth estimation
The surround view setting offers an ego-centric 360-degree perception solution [35, 39, 3, 14]. [17] introduces the surround view benchmark in the supervised setting, which learns the depth scale directly from the ground truth depth map. FSM [16] is a pioneering work in scale-aware surround view depth estimation relying on stereo constraint [53]. However, subsequent studies [44, 28] have found that reproducing the performance of [16] is challenging. Surrounddepth [44] improves the scale supervision signal using sparse point clouds from Structure-from-Motion (SFM). Building on the spatial and temporal constraints in FSM, [28] introduces a volume feature fusion module to enhance performance. For better performance, [38] proposes the temporal offline refinement strategy based on the multiple cameras and monocular depth refinement. These works [16, 44, 28, 38] all use traditional projection and index interpolation for cross-view synthesis, computing the loss between the synthesized view and target images. In contrast, our approach employs Gaussian splatting with a dedicated design, achieving better performance compared to the solely spatial constraint.
2.2 Surround view 3D occupancy estimation
Surround view 3D occupancy estimation has gained significant attention in recent years, with several benchmarks based on the Nuscene dataset [45, 41, 43]. In addition to the advanced architectures being proposed [32, 56, 51, 31], another research trend involves utilizing volume rendering for 3D supervision with 2D supervision [12, 37, 54, 24]. SimpleOcc [12] pioneered the use of volume rendering for 3D occupancy estimation, exploring both supervised and self-supervised learning. RenderOcc [37] extends semantic information for rendering. OccNeRF [54] and SelfOcc [24] share a similar approach by using 2D open-vocabulary semantic models to generate semantic maps for supervision. However, since volume rendering processes are conducted at real-world scale, these self-supervised methods [12, 54, 24] require ground truth 6D poses from sensors to provide the real-world scale for training. In contrast, we are exploring a solution that utilizes the overlap region in adjacent cameras to learn the real-world scale, eliminating the need for ground truth 6D poses.
2.3 3D Gaussian splatting
3D Gaussian splatting has become a popular method for modeling 3D and 4D scenes using well-posed images [27, 26, 8, 47], which has the property of fast rendering compared to the volume rendering in neural radiance field [36, 19, 11]. In the context of driving scenes, a line of research has focused on scene-specific reconstruction [22, 58, 49]. Our work, however, investigates the function of Gaussian splatting in a generalized setting. In this setting, existing methods generally construct 3D Gaussians from the unprojection of learned 2D Gaussian attributes [7, 57, 33, 40]. We also employ this unprojection approach, but uniquely, our approach constructs cross-view information from adjacent views to learn scale information through Gaussian splatting projection.
Additionally, two recent works, GaussianFormer [25] and GaussianBeV [6], are closely related to ours in their focus on 3D occupancy estimation and BEV prediction. However, our exploration diverges by focusing on two new properties that Gaussian splatting can contribute to occupancy estimation: scale-aware training and faster rendering.
3 Method
Figure 2: The overview of the proposed GaussianOcc. With a sequence of surround images, we employ a U-Net architecture to predict Gaussian attributes in the 2D image grid space for cross-view Gaussian splatting. This approach provides scale information in the joint training with the 6D pose network (Stage 1). For 3D occupancy estimation, we lift the 2D features to a 3D voxel space and propose voxel grid Gaussian splatting for fast rendering (Stage 2). We omit the line to from the 6D pose net to the loss in stage 1 for clarity.
3.1 Preliminaries
In this paper, we investigate the use of 3D Gaussian splatting [27] for fully self-supervised 3D occupancy estimation. 3D Gaussian Splatting (3D-GS) is a technique for modeling static 3D scenes using point primitives. Each point is represented as a scaled Gaussian distribution characterized by a 3D covariance matrix $\Sigma$ and a mean $\mu$. The Gaussian distribution for a point $\mathcal{X}$ is given by:
$G(\mathcal{X})=e^{-\frac{1}{2}(\mathcal{X}-\mu)^{T}\Sigma^{-1}(\mathcal{X}-\mu)}$ (1)
To enable efficient optimization via gradient descent, the covariance matrix $\Sigma$ is decomposed into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ as follows:
$\Sigma=\mathbf{RSS}^{T}\mathbf{R}^{T}$ (2)
The projection of Gaussians from 3D space onto a 2D image plane involves a view transformation $\mathbf{W}$ and the Jacobian of the affine approximation of the projective transformation $\mathbf{J}$. The 2D covariance matrix $\Sigma^{\prime}$ is computed as:
$\Sigma^{\prime}=\mathbf{JW\Sigma W}^{T}\mathbf{J}^{T}$ (3)
Subsequently, an alpha-blend rendering technique, similar to that used in NeRF [36], is applied. This is formulated as:
$\mathbf{C}_{\text{color}}=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i})$ (4)
Here, $c_{i}$ represents the color of each point, and the density $\alpha_{i}$ is determined by the product of a 2D Gaussian with covariance $\Sigma^{\prime}$ and a learned per-point opacity. The color is defined using spherical harmonics (SH) coefficients as described in [10, 27].
In summary, the basic 3D Gaussian Splatting approach characterizes each Gaussian point with the following attributes: (1) a 3D position $\mathcal{X}\in\mathbb{R}^{3}$, (2) a color defined by SH coefficients $c\in\mathbb{R}^{k}$ (where $k$ denotes the dimensionality of the SH basis), (3) a rotation represented by a quaternion $r\in\mathbb{R}^{4}$, (4) a scaling factor $s\in\mathbb{R}^{3}_{+}$, and (5) an opacity $\alpha\in[0,1]$.
3.2 Overview of the proposed GaussianOcc
The overview of the GaussianOcc is illustrated in Figure 2. The key contributions are the introduction of cross-view Gaussian splatting for scale-aware training and voxel grid Gaussian splatting for faster rendering. Unlike the original Gaussian splatting used for scene-specific 3D reconstruction, we explore a setting where Gaussian attributes are well-aligned in both 2D and 3D grids. This approach models the 3D scene in the 2D image plane as a depth map and in 3D grid space as a voxel (occupancy) format.
3.3 Scale-ware training by Gaussian Splatting
Scale from spatial camera rig: Similar to the previous work [16, 44], the scale information is from the surround camera rig. Specifically, the real-world scale could be obtained by leveraging camera extrinsic matrices which is to use spatial photometric loss in the overlap region between two adjacent views, i.e., warping ${I}^{i}_{t}$ to ${I}^{j}_{t}$:
$p^{i\rightarrow j}_{t}=K^{j}(T^{j})^{-1}T^{i}D_{t}^{i}(K^{i})^{-1}p^{i}_{t}$ (5)
where $K^{i},T^{i}$ are the intrinsic and extrinsic matrices of $i$th camera, $D_{t}^{i}$ is the predict depth map of $i$th camera, $p_{t}$ is the corresponding pixel during the warping. The warping operation is achieved by bilinear interpolation with the corresponding $p^{i\rightarrow j}_{t}$. However, as pointed out by [44], the mapping $p^{i\rightarrow j}_{t}$ is easy to be out of the image bounds with the weak supervision signal in such a small overlap region. Our practice also verifies this limitation. To provide a stronger supervision signal, [44] proposes to facilitate the Structure-from-Motion (SFM) to extract sparse depth information for the scale-aware training, which is time-consuming and not straightforward. Differently, we propose to use the Gaussian splatting for scale-aware projection to achieve the cross-view stereo constraint.
Gaussian Splatting for Projection (GSP): As illustrated in Figure 2, we use a depth network to predict the Gaussian attributes in 2D grid space, including the depth map, scale map, and rotation map. For each adjacent view, we first calculate the mask in the overlap region, then mask out one side of these overlap regions. Due to the presence of the other side’s overlap region, the unprojected 3D scene remains complete, which is critical for providing scale training, as indicated in the experiment section. We then perform splatting rendering on the adjacent views to obtain the rendered image. If the depth map is accurately learned, the rendered image should resemble the original images, providing the necessary scale information for the entire pipeline.
The process of acquiring the overlap mask is shown in Figure 3. The acquisition of the overlap mask is based on volume rendering. We densely sample points along the ray in one view, and a pixel is considered part of the overlap region if more than one sampled 3D point falls within the adjacent view. Note that in the DDAD dataset [14], we exclude regions with self-occlusion (such as parts of the vehicle body). Finally, we apply an erosion operation from Open CV library [2] to the mask for purification.
3.4 Fast rendering by Gaussian Splatting
Inefficient performance in volume rendering: For 2D supervision (semantic and depth maps), previous methods [12, 54, 24, 37] employed volume rendering based on dense sampling. While the final 3D voxel representation for modeling the 3D scene is much quicker than the original implicit representation (MLPs) [36], it remains time-consuming, particularly when incorporating semantic map rendering. For example, in OccNeRF [54], the number of sampled points at a resolution of $180\times 320$ is 108,735,066. However, the target optimized points correspond to the vertices in the 3D voxel grid, which total $320\times 320\times 24=2,457,600$. This redundancy in densely sampled points aids optimization with volume rendering but is highly inefficient.
Gaussian Splatting from Voxel (GSV): As analyzed above, the target optimized points are the vertices in the 3D voxel grid, prompting us to consider directly optimizing these vertices. Therefore, we investigate the use of Gaussian splatting to further accelerate the rendering process. For each vertex in the 3D voxel grid, we obtain the necessary attributes for Gaussian splatting, such as 3D position, semantic information, and opacity. Consequently, we propose treating each vertex as a 3D Gaussian. Since these vertices are arranged in the 3D voxel space, we can model the 3D scene by optimizing the related Gaussian attributes.
3.5 Loss function
We formulate the the total loss function in:
$\mathcal{L}_{total}=\mathcal{L}_{temporal}+\mathcal{L}_{cross}+\mathcal{L}_{semantic}+\mathcal{L}_{tv},$ (6)
$(\mathcal{L}_{temporal},\mathcal{L}_{cross})=\beta\frac{1-\operatorname{SSIM}\left(I_{t},\hat{I}_{t}\right)}{2}+(1-\beta)\left\|I_{t}-\hat{I}_{t}\right\|,$ (7)
where the terms $I_{t}$ and $\hat{I}_{t}$ refer to the target image and the corresponding synthesized image, respectively. Note that $\hat{I}_{t}$ in the temporal-view photometric loss, $\mathcal{L}_{temporal}$, is generated by projecting pixels from the source image using the coordinate index. In contrast, $\hat{I}_{t}$ in the cross-view photometric loss, $\mathcal{L}_{cross}$, is derived from our proposed cross-view Gaussian splatting method. $\mathcal{L}_{semantic}$ is the 2D semantic loss. $\mathcal{L}_{tv}$ is the regularization loss for the whole voxel space. The balancing parameter $\beta$ is set to 0.85 to balance the two loss terms.
Figure 3: Overlap mask in nuScenes [3] and DDAD [14].
Method
GT Occ.
GT Pose
mIoU mIoU*
$\blacksquare$ barrier
$\blacksquare$ bicycle
$\blacksquare$ bus
$\blacksquare$ car
$\blacksquare$ const. veh.
$\blacksquare$ motorcycle
$\blacksquare$ pedestrian
$\blacksquare$ traffic cone
$\blacksquare$ trailer
$\blacksquare$ truck
$\blacksquare$ drive. suf.
$\blacksquare$ sidewalk
$\blacksquare$ terrain
$\blacksquare$ manmade
$\blacksquare$ vegetation
MonoScene [4] $\times$ 6.33 6.06 7.23 4.26 4.93 9.38 5.67 3.98 3.01 5.90 4.45 7.17 14.91 7.92 7.43 1.01 7.65
BEVDet  [21] $\times$ 20.03 19.38 30.31 0.23 32.26 34.47 12.97 10.34 10.36 6.26 8.93 23.65 52.27 26.06 22.31 15.04 15.10
BEVFormer [30] $\times$ 24.64 23.67 38.79 9.98 34.41 41.09 13.24 16.50 18.15 17.83 18.66 27.70 48.95 29.08 25.38 15.41 14.46
OccFormer [56] $\times$ 22.39 21.93 30.29 12.32 34.40 39.17 14.44 16.45 17.22 9.27 13.90 26.36 50.99 34.66 22.73 6.76 6.97
RenderOcc [37] $\times$ 24.53 23.93 27.56 14.36 19.91 20.56 11.96 12.42 12.14 14.34 20.81 18.94 68.85 42.01 43.94 17.36 22.61
TPVFormer [23] $\times$ 28.69 27.83 38.90 13.67 40.78 45.90 17.23 19.99 18.85 14.30 26.69 34.17 55.65 37.55 30.70 19.40 16.78
CTF-Occ [41] $\times$ 29.54 28.53 39.33 20.56 38.29 42.24 16.93 24.52 22.72 21.05 22.98 31.11 53.33 37.98 33.23 20.79 18.00
SimpleOcc [12] $\times$ 7.99 7.05 0.67 1.18 3.21 7.63 1.02 0.26 1.80 0.26 1.07 2.81 40.44 18.30 17.01 13.42 10.84
SelfOcc [24] $\times$ 10.54 9.30 0.15 0.66 5.46 12.54 0.00 0.80 2.10 0.00 0.00 8.25 55.49 26.30 26.54 14.22 5.60
OccNeRF [54] $\times$ 10.81 9.54 0.83 0.82 5.13 12.49 3.50 0.23 3.10 1.84 0.52 3.90 52.62 20.81 24.75 18.45 13.19
GaussianOcc $\times$ $\times$ 11.26 9.94 1.79 5.82 14.58 13.55 1.30 2.82 7.95 9.76 0.56 9.61 44.59 20.10 17.58 8.61 10.29
Table 1: 3D Occupancy prediction performance on the Occ3D-nuScenes dataset in mIoU metric. Since ‘other’ and ‘other flat’ classes are the invalid prompts for open-vocabulary models, we also calculate ‘mIoU’ as the result ignoring the classes that do not consider these two classes during evaluation, while ‘mIoU*’ is the original result. GT Occ. means using the ground truth occupancy label for the supervision. GT Pose is the ground truth pose from the sensor for self-supervised geometry learning.
Method
GT Occ.
GT Pose
Backbone Input Size Epoch RayIoU RayIoU1m, 2m, 4m mIoU* FPS
[…]
Table 2: 3D Occupancy prediction performance on the Occ3D-nuScenes dataset in RayIoU metric. GT Occ. means using the ground truth occupancy label for the supervision. GT Pose is the ground truth pose from the sensor for self-supervised geometry learning. “8f” and “16f” mean fusing temporal information from 8 or 16 frames. mIoU* is the mean Intersection over Union for all categories. FPS means frame per second for each method, which is measured on a Tesla A100 GPU.
Method GT pose Occ. Abs Rel Sq Rel RMSE RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
nuScenes [3]
FSM [16] $\times$ $\times$ 0.297 - - - - - -
FSM* [16] $\times$ $\times$ 0.319 7.534 7.860 0.362 0.716 0.874 0.931
SurroundDepth [44] $\times$ $\times$ 0.280 4.401 7.467 0.364 0.661 0.844 0.917
SA-FSM [50] $\times$ $\times$ 0.272 4.706 7.391 0.355 0.689 0.868 0.929
VFF [28] $\times$ $\times$ 0.289 5.718 7.551 0.348 0.709 0.876 0.932
R3D3 [38] $\times$ $\times$ 0.253 4.759 7.150 - 0.729 - -
GaussianOcc $\ddagger$ $\times$ $\times$ 0.258 5.733 7.222 0.343 0.753 0.888 0.934
SimpleOcc [12] 0.224 3.383 7.165 0.333 0.753 0.877 0.930
OccNeRF [54] 0.202 2.883 6.697 0.319 0.768 0.882 0.931
SelfOcc [24] 0.215 2.743 6.706 0.316 0.753 0.875 0.932
GaussianOcc $\times$ 0.211 3.115 7.131 0.326 0.762 0.878 0.931
DDAD [14]
FSM*  [16] $\times$ $\times$ 0.228 4.409 13.433 0.342 0.687 0.870 0.932
VFF [28] $\times$ $\times$ 0.218 3.660 13.327 0.339 0.674 0.862 0.932
SurroundDepth [44] $\times$ $\times$ 0.208 3.371 12.977 0.330 0.693 0.871 0.934
SA-FSM [50] $\times$ $\times$ 0.187 3.093 12.578 0.311 0.731 0.891 0.945
R3D3 [38] $\times$ $\times$ 0.162 3.019 11.408 - 0.811 - -
GaussianOcc $\ddagger$ $\times$ $\times$ 0.212 3.556 12.564 0.320 0.701 0.888 0.944
GaussianOcc $\times$ 0.228 3.854 14.326 0.357 0.660 0.853 0.922
Table 3: Comparisons for self-supervised multi-camera depth estimation on the nuScenes  [3] and DDAD datasets  [14]. The results are averaged over all views without median scaling at test time. ‘FSM*’ is the reproduced result in [28]. GaussianOcc $\ddagger$ represents the depth estimation result from Stage 1. Occ. represents the ability of the method to predict the 3D occupancy.
Figure 4: The visualization of the render depth map and 3D occupancy prediction on nuScenes and DDAD dataset.
4 Experiment
4.1 Tasks, Datasets, and Metric
nuScenes: We perform 3D occupancy estimation and depth estimation using the nuScenes dataset [3]. For 3D occupancy estimation, we utilize annotations from Occ3D-nuScenes [41], which includes 600 scenes for training and 150 scenes for evaluation. Each sample has a perception range of [-40m, -40m, -1m, 40m, 40m, 5.4m] with a voxel size of 0.4m. For fair comparison, we use the 2D semantic map provided by OccNeRF [54]. We measure 3D occupancy estimation performance using the mean Intersection over Union (mIoU) metric. For depth estimation, we set the perception range in [-80m, -80m, -1m, 80m, 80m, 6m], while we clamp the ground truth to a range of 0.1m to 80m for evaluation, consistent with OccNeRF and SurroundDepth [44]. We evaluate depth maps using error metrics (Abs Rel, Sq Rel, RMSE, RMSE log) and threshold accuracy metrics ($\delta$).
DDAD: We conduct 3D occupancy estimation and depth estimation using the DDAD dataset [14]. Due to the lack of ground truth labels, we only present qualitative results for the fully self-supervised 3D occupancy estimation setting. Following the procedure used with the nuScenes dataset, we obtain the 2D semantic labels as per the OccNeRF pipeline. For depth estimation, we also set the perception range in [-80m, -80m, -1m, 80m, 80m, 6m], but clamp the ground truth to a range of 0.1m to 200m for evaluation, consistent with SurroundDepth [44].
Scale aware training in [44] Scale aware training by ours RMSE
Spatial loss SFM loss GS loss Mask Erode Refine Abs Rel Sq Rel RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
✓* 0.970 13.702 17.931 3.654 0.000 0.000 0.001
0.921 30.393 14.444 0.808 0.057 0.187 0.667
✓* ✓* 0.280 4.401 7.467 0.364 0.661 0.844 0.917
0.672 25.405 11.999 0.568 0.419 0.808 0.878
0.798 11.571 15.251 1.472 0.006 0.015 0.028
0.293 7.127 7.536 0.376 0.743 0.876 0.923
0.281 6.986 7.347 0.354 0.766 0.885 0.929
0.258 5.733 7.222 0.343 0.753 0.888 0.934
Table 4: Ablation study for scale-aware depth estimation on the Nuscene dataset [15]. The scores are averaged over all views without median-scaling at test time. ✓* means the result from the original paper and ✓means the result from our implementation in New-CRFs network [52]. Spatial loss denotes the spatial context loss used in [44] and SFM loss denotes the sparse depth pretrain from the off-line Structure-from-Motion process. GS loss means using the spatial context constraint by our proposed Gaussian splatting for projection. Mask represents using the mask-out strategy before the unprojection. Erode means use the erode process to the binary overlap mask and Refine is the refinement of the depth estimation network by fixing the 6D pose net.
4.2 Implementation details
Network details: For U-Net architecture, we adopt the New-CRFs [52] as the depth estimation network, which is based on the Swin Transformer [34]. The 6D pose network is the same as that used in SurroundDepth [44]. For the 2D to 3D lifting, we follow the approach used in SimpleOcc [12]. In depth estimation benchmark, we use the network proposed by SimpleOcc, where the final output size is 256×256×16. In our Gaussian splatting setting, we further upsample the final output to 512×512×32 for improved performance, which only requires minimal additional computational cost. For occupancy estimation, we use the same network as OccNeRF [54] to ensure a fair comparison.
Training details: We propose a two-stage training process for fully self-supervised 3D occupancy estimation. In the first stage, we jointly train the depth estimation network and the 6D pose network. In the second stage, we train the 3D occupancy network without optimizing the 6D pose network. During the first stage, we train the models for 8 epochs on the nuScenes dataset and 12 epochs on the DDAD dataset. In the second stage, we train the models for 12 epochs on both the nuScenes and DDAD datasets. The optimizer and learning rate adjustment strategy are the same as those used in SimpleOcc [12] and OccNeRF [54].
4.3 Main results
3D occupancy estimation in mIoU metric: As shown in Table 1, the proposed GaussianOcc achieves the best performance compared with other self-supervised methods, without the need for ground truth occupancy labels and ground truth 6D poses for training. It’s worth noting that SelfOcc [24] differs slightly from methods like SimpleOcc, OccNeRF, and our approach. Unlike these methods, SelfOcc does not predict semantic information directly in the 3D occupancy space. Instead, it attaches 2D semantic information to the 3D occupied voxels, where the 2D semantic data is sourced from a third-party open-vocabulary model [55].
3D occupancy estimation in RayIoU metric: We also compare our method using a newly proposed metric, RayIoU, introduced by the recent work [32]. The RayIoU is a ray-based evaluation metric to solve the inconsistency penalty along the depth axis raised in traditional voxel-level mIoU criteria. As shown in Table 2, our approach also outperforms OccNeRF [54]. It’s important to note that the FPS is calculated excluding rendering time. Since GaussianOcc and OccNeRF utilize the same network architecture, they share the same inference time when the rendering process is not taken into account.
3D occupancy estimation in DDAD: Thanks to our fully self-supervised learning approach, we are able to perform 3D occupancy estimation on a similar dataset, DDAD [14]. To the best of our knowledge, our work is the first to introduce 3D occupancy estimation on this dataset, although we cannot provide quantitative results due to the lack of 3D annotations. We present the visualization results in Figure 4 and more results could be found in the supplement material.
Depth estimation: We present a comparison of depth estimation results in Table 3 for both the nuScenes and DDAD datasets. In the first stage of monocular depth estimation, GaussianOcc $\ddagger$ achieves top performance on the nuScenes dataset and delivers competitive results on the DDAD dataset. It is important to note that methods such as SurroundDepth [44] and SA-FSM [50] rely on third-party sparse point cloud extraction. Additionally, R3D3 [38] is a temporal offline refinement method that requires multi-frame optimization, as discussed in [24]. In the second stage, which involves depth estimation from rendering, our method also achieves competitive results compared to those trained with ground truth poses. Besides, we observe that the rendered depth in stage two outperforms the depth results from stage 1 on the nuScenes dataset, whereas the opposite is true for the DDAD dataset. This discrepancy might be attributed to differences in perception range—80 meters in nuScenes versus 200 meters in DDAD.
4.4 Ablation Study
Scale aware training: To demonstrate the effectiveness of the proposed scale-aware training using Gaussian Splatting, we conducted experiments comparing our method with existing approaches that use SFM loss for training [44]. As indicated in Table 4, we implemented scale-aware training from [44] on our network [52], but the results were suboptimal, and our proposed method outperformed the results reported in the original paper. In our method, we observed that a naive implementation of Gaussian splatting without a masking strategy is ineffective because it lacks cross-view rendering, which is essential for applying cross-view constraints. To enhance performance, we introduced an erosion operation on the binary mask to purify it, excluding regions that may fall outside the overlap area. This step ensures better alignment during training. Finally, we refined the depth estimation by fixing the 6D pose network and disabling the cross-view loss. This refinement helps reduce artifacts at the edges of the overlap region. We present the cross-view splatting render result in Figure 5, where the rendering result in overlap regions is reasonable in the correct depth estimation from cross view.
6D pose and training strategy: Building on the success of scale-aware training, we compared the depth estimation performance of our proposed two-stage training method with that of models trained using ground truth (GT) pose. As shown in Table 5, our two-stage approach delivers competitive results when compared to training with GT pose, indicating that the predicted pose is of high quality. Additionally, the GT pose provided by the nuScenes dataset is limited by sensor inaccuracies, particularly a lack of vertical movement. In contrast, the learned 6D pose in our method does not suffer from this limitation. Furthermore, in our one-stage experiment, we found that this approach was not effective. The cross-view loss in the 3D voxel space led to local optimization, which failed to generalize predictions to non-overlapping regions as indicated in Figure 6. However, this issue did not arise in the depth maps generated by the 2D decoder in stage one.
Volume rendering and Splatting rendering: In Table 6, we examine the performance of volume rendering (VR) and splatting rendering (SR) in the voxel space. For the scale of each 3D Gaussian, we set a uniform scale $s$ for all Gaussians, considering the well-arranged positions of vertices within the voxel grid. Since the scale $s$ is consistent across all three dimensions, we set the rotation $\mathbf{R}$ as the identity matrix. We first explored the optimal scale and found that a scale of 0.1 produced the best results. Notably, the performance of VR and SR was very close when using this scale. The render depth maps are shown in Figure 6. The depth map by the volume rendering is smoother thanks to dense sampling and the depth maps with splatting tendering have the graininess effect, especially at the small scale factor.
Pose and render types ablation study in occupancy task: In Table 7, we conduct an ablation study to evaluate the impact of using ground truth (GT) versus learned poses in the occupancy task, with training conducted using different rendering methods. The results show that splatting rendering achieves superior performance in both occupancy metrics (mIoU) and depth metrics. One observation is that depth estimation results from 3D semantic learning with volume rendering are significantly worse than those in Table 3 and the proposed splatting rendering method maintains consistent performance, which indicates the splatting rendering in voxel grid may contribute better 3D geometry.
Rendering efficiency analysis: Finally, we analyze rendering efficiency in Table 8. The results indicate that volume rendering consumes 5 times more rendering time compared to splatting rendering, and as the resolution increases, both rendering time and GPU consumption rise significantly. In contrast, splatting rendering shows no significant increase in computational cost with higher resolutions, highlighting its efficiency and scalability.
Pose type Abs Rel Sq Rel RMSE $\delta<1.25$
GT pose 0.214 3.362 7.127 0.771
Two stages 0.211 3.115 7.131 0.762
One stage 0.946 17.008 16.397 0.103
Table 5: Comparison of the ground truth pose (GT pose) and our learned pose (Two stages) on the depth estimation task [15]. We also conducted the experiment that directly uses the cross-view loss to render a depth map (One stage).
Render type Abs Rel Sq Rel RMSE $\delta<1.25$
VR 0.215 3.508 7.113 0.775
SR ($s$ = 0.05) 0.223 3.694 7.246 0.761
SR ($s$ = 0.1) 0.217 3.504 7.152 0.770
SR ($s$ = 0.15) 0.217 3.406 7.204 0.763
Table 6: Comparison of the render result between the volume rendering (VR) [54] and splatting rendering (SR, Ours) on depth estimation task [15]. The coefficient $s$ means the scale of the 3D Gaussian. We use GT pose for this set of the ablation study.
Pose type mIoU Abs Rel Sq Rel $\delta<1.25$
GT pose (VR) 10.81 0.456 12.682 0.704
GT pose (SR) 11.30 0.225 4.339 0.787
Learned pose (VR) 11.19 0.506 15.577 0.684
Learned pose (SR) 11.43 0.219 3.625 0.778
Table 7: Comparison of ground truth pose (GT Pose) and our learned pose (Two stages) on 3D occupancy estimation task [15] in volume rendering (VR) and splatting rendering (SR).
Render Render resolution and time (s) Training
type 180 $\times$ 320 240 $\times$ 520 360 $\times$ 640 time (h)
VR $\approx$ 0.85 $\approx$ 1.57 N/A $\approx$ 2.68
SR $\approx$ 0.17 $\approx$ 0.17 $\approx$ 0.17 $\approx$ 1
Table 8: Comparison of rendering efficiency between volume rendering (VR) [54] and splatting rendering (SR, Ours) on 3D occupancy estimation task [15]. The render time is calculated from surround 6 images. ”N/A” indicates out-of-memory errors running in NVIDIA A 100 (40 GB). Training time is averaged per epoch.
Figure 5: The visualization for the depth map, splatting render image and depth map from cross view in stage 1.
Figure 6: The comparison for the depth map in the different setting, corresponding to the training strategy in Table 5 and rendering type in Table 6.
5 Conclusion
In this paper, we introduce GaussianOcc, a fully self-supervised and efficient approach for 3D occupancy estimation. Through the carefully designed cross-view splatting rendering, we can accurately learn the real scale in depth and the 6D pose net, enabling effective self-supervised 3D occupancy learning. Additionally, the proposed Gaussian splatting in voxel grids outperforms volume rendering in 3D occupancy estimation while reducing computational cost. We validate our method on the nuScenes and DDAD datasets, demonstrating its strong generalization ability to surround view scenes.
Appendix
Appendix A More implementation details
The detailed parameter setting in depth estimation network [52]. During the joint depth and 6D pose training in stage 1, we predict the 3D Gaussian parameters alongside the 2D depth map. Since the Gaussian parameters are well-arranged in the 2D image plane prior to unprojection, we maintain equal scaling across all three dimensions of each 3D Gaussian and constrain the maximum scale to 0.02. Given that the scale $s$ is uniform across all dimensions, we set the rotation matrix $\mathbf{R}$ to the identity matrix. Additionally, we assign an opacity value of 1 to each 3D Gaussian, ensuring that every 2D depth value corresponds to a valid point in 3D space.
The detailed parameter setting in voxel grid splatting rendering. We chose a fixed scale for each grid vertex to ensure a well-arranged structure that accurately models the 3D space. If we were to use a learnable scale, it could lead to a situation where the scale is small but the opacity is large, which might not be captured in the rendered depth map and semantic map but could still impact the 3D occupancy result. Therefore, using a fixed scale is sufficient for optimization, as demonstrated in the results presented in the main paper.
[…]
Detailed definition of depth map metric. Following the depth estimation task [44], we report the depth map evaluation with the following metrics,
[…]
where $M$ is the valid pixel, $\hat{d}$ is the ground truth depth and $d^{*}$ is the predicted depth.
Appendix B More experiment results
More visualization. We provide more visualization for nuScenes dataset in Figure A and DDAD dataset in Figure B. Please check the attached videos for sequence visualization.
Appendix C Limitation and future work
The proposed method achieves reasonable predictions in most scenes; however, we observe that some cases still present challenges, as shown in Figure C. Specifically, in the DDAD dataset, incorrect predictions occur in the back camera in certain situations as marked with the red circle, where the drivable surface is mistakenly projected into the car due to extensive self-occlusion. Notably, this issue is absent in the nuScenes dataset, which has less self-occlusion. We believe that this problem could be mitigated with better 2D semantic maps for supervision, which warrants further investigation. Additionally, we plan to explore the use of splatting rendering for 3D occupancy flow in future work.
Figure A: The visualization of the render depth map and 3D occupancy prediction on nuScenes dataset.
Figure B: The visualization of the render depth map and 3D occupancy prediction on DDAD dataset
Figure C: Some wrong predictions due to the large self-occlusion on DDAD dataset.

### v3

@@
 Abstract
-We introduce GaussianOcc, a systematic method that investigates the two usages of Gaussian splatting for fully self-supervised and efficient 3D occupancy estimation in surround views.
+We introduce GaussianOcc, a systematic method that investigates two usages of Gaussian splatting for fully self-supervised and efficient 3D occupancy estimation in surround views.
 First, traditional methods for self-supervised 3D occupancy estimation still require ground truth 6D poses from sensors during training.
 To address this limitation, we propose Gaussian Splatting for Projection (GSP) module to provide accurate scale information for fully self-supervised training from adjacent view projection.
-Additionally, existing methods rely on volume rendering for final 3D voxel representation learning using 2D signals (depth maps, semantic maps), which is both time-consuming and less effective.
+Additionally, existing methods rely on volume rendering for final 3D voxel representation learning using 2D signals (depth maps and semantic maps), which is time-consuming and less effective.
 We propose Gaussian Splatting from Voxel space (GSV) to leverage the fast rendering properties of Gaussian splatting.
-As a result, the proposed GaussianOcc method enables fully self-supervised (no ground truth pose) 3D occupancy estimation in competitive performance with low computational cost (2.7 times faster in training and 5 times faster in rendering).
-The relevant code will be available in https://github.com/GANWANSHUI/GaussianOcc.git.
+As a result, the proposed GaussianOcc method enables fully self-supervised (no ground-truth pose) 3D occupancy estimation in competitive performance with low computational cost (2.7 times faster in training and 5 times faster in rendering).
+The relevant code is available in https://github.com/GANWANSHUI/GaussianOcc.git.
+Abstract
+In this supplementary material, we provide more implementation details, experiment results, and further discussion on the limitations and future work.
+1 Introduction
 Figure 1: Problem setting of GaussianOcc.
-Given the surround image sequence and its corresponding 2D semantic annotation, GaussianOcc is able to perform 3D occupancy estimation without the need for the ground truth occupancy label and ground truth 6D pose for training.
-1 Introduction
+Given a surround image sequence and its corresponding 2D semantic annotation, GaussianOcc is able to perform 3D occupancy estimation without the need for ground truth occupancy label and ground truth 6D pose for training.
 In recent years, both vision-centric and grid-centric perception approaches have gained significant attention in the autonomous driving industry and academia [34, 39].
@@
 Volume rendering allows 3D representation learning using 2D supervision signals, such as 2D semantic maps and depth maps, thereby eliminating the need for extensive 3D annotation during training.
-The existing methods [54, 24] achieve self-supervised learning through volume rendering, where the 2D semantic map supervision is derived from open-vocabulary semantic segmentation [55], and the depth map supervision is obtained from self-supervised depth estimation [13].
+Existing methods [54, 23] achieve self-supervised learning through volume rendering, where the 2D semantic map supervision is derived from open-vocabulary semantic segmentation [55], and the depth map supervision is obtained from self-supervised depth estimation [12].
 However, these approaches face two significant limitations.
@@
 These limitations impede the development of a more general and efficient paradigm for self-supervised 3D occupancy estimation.
-To address the aforementioned limitations, we explore a fully self-supervised and efficient approach to 3D occupancy estimation using Gaussian splatting [27, 1].
+To address the aforementioned limitations, we propose a fully self-supervised and efficient approach to 3D occupancy estimation using Gaussian splatting [26, 1].
 Specifically, we introduce the use of Gaussian splatting to perform cross-view splatting, where the rendered image constructs a cross-view loss that provides scale information during joint training with the 6D pose network.
-This eliminates the need for a ground truth 6D pose during training.
+This eliminates the need for ground truth 6D pose during training.
 To improve rendering efficiency, we move away from the dense sampling required in traditional volume rendering.
 Instead, we propose performing Gaussian splatting directly from the 3D voxel space.
-In this approach, each vertex in the voxel grid is treated as a 3D Gaussian, and we directly optimize the attributes of these Gaussians—such as semantic features and opacity—within the voxel space.
+In this approach, each vertex in the voxel grid is treated as a 3D Gaussian, and we optimize the attributes of these Gaussians—such as semantic features and opacity—directly within the voxel space.
 Through this novel approach, our proposed method, GaussianOcc, makes a progress toward fully self-supervised and efficient 3D occupancy estimation, as outlined in Figure 1.
-The main contributions of this work are summarized as follows.
-  • We introduce the first fully self-supervised method for efficient surrounding-view 3D occupancy estimation, featuring the exploration of the Gaussian splatting.
-  • We propose Gaussian splatting for cross-view projection, which could provide the scale information to get rid of the need of a ground truth 6D pose during the training.
-  • We propose Gaussian splatting from voxel space achieving competitive performance with 2.7 times faster training and 5 times faster rendering compared with the volume rendering.
+In summary, our core contributions are as follows:
+  • We introduce the first fully self-supervised method for efficient surrounding-view 3D occupancy estimation, featuring the exploration of Gaussian splatting.
+  • We propose Gaussian splatting for cross-view projection module, which can provide scale information to get rid of the need of ground truth 6D pose during the training.
+  • We propose Gaussian splatting from voxel space module, achieving competitive performance with 2.7 times faster training and 5 times faster rendering compared to the previous works with volume rendering.
 2 Related work
@@
 These works [15, 44, 27, 38] all use traditional projection and index interpolation for cross-view synthesis, computing the loss between the synthesized view and target images.
-In contrast, our approach employs Gaussian splatting with a dedicated design, achieving better performance compared to the solely spatial constraint.
+Compared to the traditional projection, our approach employs Gaussian splatting with a dedicated design for cross-view constraint, achieving better performance.
 2.2 Surround view 3D occupancy estimation
-Surround view 3D occupancy estimation has gained significant attention in recent years, with several benchmarks based on the Nuscene dataset [45, 41, 43].
-In addition to the advanced architectures being proposed [32, 56, 51, 31], another research trend involves utilizing volume rendering for 3D supervision with 2D supervision [12, 37, 54, 24].
+Surround view 3D occupancy estimation has gained significant attention in recent years, with several benchmarks based on the nuScenes dataset [45, 41, 43].
+In addition to the advanced architectures being proposed [31, 56, 51, 30], another research trend involves utilizing volume rendering for 3D occupancy learning with 2D supervision [11, 37, 54, 23].
 SimpleOcc [11] pioneered the use of volume rendering for 3D occupancy estimation, exploring both supervised and self-supervised learning.
@@
 However, since volume rendering processes are conducted at real-world scale, these self-supervised methods [11, 54, 23] require ground truth 6D poses from sensors to provide the real-world scale for training.
-In contrast, we are exploring a solution that utilizes the overlap region in adjacent cameras to learn the real-world scale, eliminating the need for ground truth 6D poses.
+Differently, we are exploring a solution that utilizes the overlap region in adjacent cameras to learn the real-world scale, eliminating the need for ground truth 6D poses.
+Figure 2: GaussianOcc is a two-stage method.
+In Stage 1, we train a scale-aware 6D pose network, using a U-Net architecture to predict Gaussian attributes in the 2D image grid space for cross-view Gaussian splatting.
+This approach provides scale information in the joint training with the 6D pose network.
+Based on the 6D pose from Stage 1, we perform self-supervised 3D occupancy estimation in Stage 2, where we lift the 2D features to a 3D voxel space and propose voxel grid Gaussian splatting for fast rendering.
+Note that, for clarity, we omit the line from the 6D pose network to the loss in Stage 1, and the 2D encoder is independent for each stage (not shared).
 2.3 3D Gaussian splatting
 3D Gaussian splatting has become a popular method for modeling 3D and 4D scenes using well-posed images [26, 25, 8, 47], which has the property of fast rendering compared to the volume rendering in neural radiance field [35, 18, 10].
-In the context of driving scenes, a line of research has focused on scene-specific reconstruction [22, 58, 49].
-Our work, however, investigates the function of Gaussian splatting in a generalized setting.
-In this setting, existing methods generally construct 3D Gaussians from the unprojection of learned 2D Gaussian attributes [7, 57, 33, 40].
+In driving scenes, a line of research has focused on scene-specific reconstruction [21, 58, 49].
+Our work, however, investigates the function of Gaussian splatting in a generalized setting, where existing methods generally construct 3D Gaussians from the unprojection of learned 2D Gaussian attributes [7, 57, 32, 40].
 We also employ this unprojection approach, but uniquely, our approach constructs cross-view information from adjacent views to learn scale information through Gaussian splatting projection.
-Additionally, two recent works, GaussianFormer [25] and GaussianBeV [6], are closely related to ours in their focus on 3D occupancy estimation and BEV prediction.
+In addition, we investigate the Gaussian splatting on the voxel space for faster rendering compared with the previous volume rendering-based works [11, 54, 23].
+Note that, two recent works, GaussianFormer [24] and GaussianBeV [6], are closely related to ours in their focus on 3D occupancy estimation and BEV prediction.
 However, our exploration diverges by focusing on two new properties that Gaussian splatting can contribute to occupancy estimation: scale-aware training and faster rendering.
 3 Method
-Figure 2: The overview of the proposed GaussianOcc.
-With a sequence of surround images, we employ a U-Net architecture to predict Gaussian attributes in the 2D image grid space for cross-view Gaussian splatting.
-This approach provides scale information in the joint training with the 6D pose network (Stage 1).
-For 3D occupancy estimation, we lift the 2D features to a 3D voxel space and propose voxel grid Gaussian splatting for fast rendering (Stage 2).
-We omit the line to from the 6D pose net to the loss in stage 1 for clarity.
 3.1 Preliminaries
-In this paper, we investigate the use of 3D Gaussian splatting [27] for fully self-supervised 3D occupancy estimation.
-3D Gaussian Splatting (3D-GS) is a technique for modeling static 3D scenes using point primitives.
-Each point is represented as a scaled Gaussian distribution characterized by a 3D covariance matrix $\Sigma$ and a mean $\mu$.
-The Gaussian distribution for a point $\mathcal{X}$ is given by:
-$G(\mathcal{X})=e^{-\frac{1}{2}(\mathcal{X}-\mu)^{T}\Sigma^{-1}(\mathcal{X}-\mu)}$ (1)
-To enable efficient optimization via gradient descent, the covariance matrix $\Sigma$ is decomposed into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ as follows:
-$\Sigma=\mathbf{RSS}^{T}\mathbf{R}^{T}$ (2)
-The projection of Gaussians from 3D space onto a 2D image plane involves a view transformation $\mathbf{W}$ and the Jacobian of the affine approximation of the projective transformation $\mathbf{J}$.
-The 2D covariance matrix $\Sigma^{\prime}$ is computed as:
-$\Sigma^{\prime}=\mathbf{JW\Sigma W}^{T}\mathbf{J}^{T}$ (3)
-Subsequently, an alpha-blend rendering technique, similar to that used in NeRF [36], is applied. This is formulated as:
-$\mathbf{C}_{\text{color}}=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i})$ (4)
-Here, $c_{i}$ represents the color of each point, and the density $\alpha_{i}$ is determined by the product of a 2D Gaussian with covariance $\Sigma^{\prime}$ and a learned per-point opacity.
-The color is defined using spherical harmonics (SH) coefficients as described in [10, 27].
-In summary, the basic 3D Gaussian Splatting approach characterizes each Gaussian point with the following attributes: (1) a 3D position $\mathcal{X}\in\mathbb{R}^{3}$, (2) a color defined by SH coefficients $c\in\mathbb{R}^{k}$ (where $k$ denotes the dimensionality of the SH basis), (3) a rotation represented by a quaternion $r\in\mathbb{R}^{4}$, (4) a scaling factor $s\in\mathbb{R}^{3}_{+}$, and (5) an opacity $\alpha\in[0,1]$.
-3.2 Overview of the proposed GaussianOcc
-The overview of the GaussianOcc is illustrated in Figure 2.
-The key contributions are the introduction of cross-view Gaussian splatting for scale-aware training and voxel grid Gaussian splatting for faster rendering.
-Unlike the original Gaussian splatting used for scene-specific 3D reconstruction, we explore a setting where Gaussian attributes are well-aligned in both 2D and 3D grids.
-This approach models the 3D scene in the 2D image plane as a depth map and in 3D grid space as a voxel (occupancy) format.
+3D Gaussian splatting [26] is for modeling static 3D scenes using point primitives, where each primitive is characterized with the following attributes: (1) a 3D position $\mathcal{X}\in\mathbb{R}^{3}$, (2) a color defined by SH coefficients $c\in\mathbb{R}^{k}$ (where $k$ denotes the dimensionality of the SH basis), (3) a rotation represented by a quaternion $r\in\mathbb{R}^{4}$, (4) a scaling factor $s\in\mathbb{R}^{3}_{+}$, and (5) an opacity $\alpha\in[0,1]$.
+The original Gaussian splatting is for scene-specific, fast 3D novel view synthesis, where the attributes of Gaussian points are optimized by the multi-view constraint.
+In contrast, in this work, we explore a setting where Gaussian attributes are well-aligned in both 2D and 3D grids, allowing us to benefit from the Gaussian splatting rendering for the scale-aware training by cross-view constraint and faster rendering on voxel grids as illustrated in Figure 2.
 3.2 Scale-ware training by Gaussian Splatting
-Scale from spatial camera rig: Similar to the previous work [16, 44], the scale information is from the surround camera rig.
-Specifically, the real-world scale could be obtained by leveraging camera extrinsic matrices which is to use spatial photometric loss in the overlap region between two adjacent views, i.e., warping ${I}^{i}_{t}$ to ${I}^{j}_{t}$:
+Scale from spatial camera rig: Similar to the previous work [15, 44, 27], the scale information is from the surround camera rig.
+Specifically, the real-world scale can be obtained by leveraging camera extrinsic matrices which is to use spatial photometric loss in the overlap region between two adjacent views, i.e., warping ${I}^{i}_{t}$ to ${I}^{j}_{t}$:
 $p^{i\rightarrow j}_{t}=K^{j}(T^{j})^{-1}T^{i}D_{t}^{i}(K^{i})^{-1}p^{i}_{t}$ (1)
-where $K^{i},T^{i}$ are the intrinsic and extrinsic matrices of $i$th camera, $D_{t}^{i}$ is the predict depth map of $i$th camera, $p_{t}$ is the corresponding pixel during the warping.
+where $K^{i},T^{i}$ are the intrinsic and extrinsic matrices of $i$-th camera, $D_{t}^{i}$ is the predicted depth map of $i$-th camera, $p_{t}$ is the corresponding pixel during the warping.
 The warping operation is achieved by bilinear interpolation with the corresponding $p^{i\rightarrow j}_{t}$.
-However, as pointed out by [44], the mapping $p^{i\rightarrow j}_{t}$ is easy to be out of the image bounds with the weak supervision signal in such a small overlap region.
-Our practice also verifies this limitation.
-To provide a stronger supervision signal, [44] proposes to facilitate the Structure-from-Motion (SFM) to extract sparse depth information for the scale-aware training, which is time-consuming and not straightforward.
-Differently, we propose to use the Gaussian splatting for scale-aware projection to achieve the cross-view stereo constraint.
-Gaussian Splatting for Projection (GSP): As illustrated in Figure 2, we use a depth network to predict the Gaussian attributes in 2D grid space, including the depth map, scale map, and rotation map.
+However, as pointed out by [44], the mapping $p^{i\rightarrow j}_{t}$ can easily go out of image bounds with the weak supervision signal in such a small overlap region.
+Therefore, apart from the spatial loss, [44] proposes to facilitate the Structure-from-Motion (SFM) to extract sparse depth information for the direct depth supervision to provide a stronger supervision signal, but it is time-consuming and not straightforward.
+Different from [44], [27] enhances the depth estimation performance with spatio-temporal context that does not need the sparse depth from SFM but the performance is still limited.
+Inspired by the explicit sparse depth supervision in [44], we ask whether we can enforce the cross-view constraint on adjacent views more explicitly.
+The answer is yes.
+We find that the nature of Gaussian splatting is scale-aware projection that could serve for the cross-view stereo constraint.
+We propose Gaussian splatting for projection in stage 1 for better scale-aware training.
+Gaussian Splatting for Projection (GSP): As illustrated in Figure 2, we adapt a depth network [52] to predict the Gaussian attributes in 2D grid space, where, apart from the original depth map, we also predict the scale map and rotation map.
 For each adjacent view, we first calculate the mask in the overlap region, then mask out one side of these overlap regions.
-Due to the presence of the other side’s overlap region, the unprojected 3D scene remains complete, which is critical for providing scale training, as indicated in the experiment section.
+Due to the presence of the other side’s overlap region, the unprojected 3D scene remains complete, which is critical for providing scale training, as indicated in the experiment section (Table 3).
 We then perform splatting rendering on the adjacent views to obtain the rendered image.
 If the depth map is accurately learned, the rendered image should resemble the original images, providing the necessary scale information for the entire pipeline.
-The process of acquiring the overlap mask is shown in Figure 3.
-The acquisition of the overlap mask is based on volume rendering.
+The process of acquiring the overlap mask is illustrated in Figure 3.
+The acquisition of the overlap mask is similar to volume rendering.
 We densely sample points along the ray in one view, and a pixel is considered part of the overlap region if more than one sampled 3D point falls within the adjacent view.
+The overlap mask is determined by the camera extrinsic and defined max depth (e.g., 80m in nuScenes).
 Note that in the DDAD dataset [13], we exclude regions with self-occlusion (such as parts of the vehicle body).
@@
 3.3 Fast rendering by Gaussian Splatting
-Inefficient performance in volume rendering: For 2D supervision (semantic and depth maps), previous methods [12, 54, 24, 37] employed volume rendering based on dense sampling.
-While the final 3D voxel representation for modeling the 3D scene is much quicker than the original implicit representation (MLPs) [36], it remains time-consuming, particularly when incorporating semantic map rendering.
+Inefficient performance in volume rendering: For 2D supervision (semantic and depth maps), previous methods [11, 54, 23, 37] employed volume rendering based on dense sampling.Although the final 3D voxel representation for modeling the 3D scene is much quicker than the original implicit representation [35], it remains time-consuming, particularly when incorporating semantic map rendering.
 For example, in OccNeRF [54], the number of sampled points at a resolution of $180\times 320$ is 108,735,066.
-However, the target optimized points correspond to the vertices in the 3D voxel grid, which total $320\times 320\times 24=2,457,600$.
-This redundancy in densely sampled points aids optimization with volume rendering but is highly inefficient.
+However, the target optimized points correspond to the vertices in the 3D voxel grid, which total $320\times 320\times 24=2,577,525$.
+This redundancy in densely sampled points helps optimization with volume rendering but is highly inefficient.
 Gaussian Splatting from Voxel (GSV): As analyzed above, the target optimized points are the vertices in the 3D voxel grid, prompting us to consider directly optimizing these vertices.
-Therefore, we investigate the use of Gaussian splatting to further accelerate the rendering process.
-For each vertex in the 3D voxel grid, we obtain the necessary attributes for Gaussian splatting, such as 3D position, semantic information, and opacity.
-Consequently, we propose treating each vertex as a 3D Gaussian.
-Since these vertices are arranged in the 3D voxel space, we can model the 3D scene by optimizing the related Gaussian attributes.
+Interestingly, we find that it is suitable to use the Gaussian splatting to replace the volume rendering if we regard the vertices of each the voxel grid as the position of the 3D Gaussian.
+Then, we can optimize the attributes of the 3D Gaussian, such as semantic and opacity information.
+For example, in the empty space of the voxel, even though we have the vertices at that region during the splatting rendering, after the optimization, the network would predict the opacity as zero at these vertices then these vertices would not contribute any geometry or semantic information during the rendering.
+Since all the vertices are arranged in 3D voxel space with real-world 3D position $\mathcal{X}$, we use fixed scale $s$ and fixed rotation $r$ for each vertex for simplification.
+This allows us to model the 3D scene by optimizing the rest Gaussian attributes (semantic and opacity).
 3.4 Loss function
-We formulate the the total loss function in:
-$\mathcal{L}_{total}=\mathcal{L}_{temporal}+\mathcal{L}_{cross}+\mathcal{L}_{semantic}+\mathcal{L}_{tv},$ (6)
-$(\mathcal{L}_{temporal},\mathcal{L}_{cross})=\beta\frac{1-\operatorname{SSIM}\left(I_{t},\hat{I}_{t}\right)}{2}+(1-\beta)\left\|I_{t}-\hat{I}_{t}\right\|,$ (7)
-where the terms $I_{t}$ and $\hat{I}_{t}$ refer to the target image and the corresponding synthesized image, respectively.
-Note that $\hat{I}_{t}$ in the temporal-view photometric loss, $\mathcal{L}_{temporal}$, is generated by projecting pixels from the source image using the coordinate index.
-In contrast, $\hat{I}_{t}$ in the cross-view photometric loss, $\mathcal{L}_{cross}$, is derived from our proposed cross-view Gaussian splatting method.
-$\mathcal{L}_{semantic}$ is the 2D semantic loss.
-$\mathcal{L}_{tv}$ is the regularization loss for the whole voxel space.
-The balancing parameter $\beta$ is set to 0.85 to balance the two loss terms.
+We formulate the loss function of each stage as follows:
+$\mathcal{L}_{stage1}=\mathcal{L}_{temporal}+\mathcal{L}_{cross},$ (2)
+$\mathcal{L}_{stage2}=\mathcal{L}_{temporal}+\mathcal{L}_{semantic},$ (3)
+$(\mathcal{L}_{temporal},\mathcal{L}_{cross})=\frac{1-\operatorname{SSIM}\left(I_{t},\hat{I}_{t}\right)}{2}+(\beta)\left\|I_{t}-\hat{I}_{t}\right\|,$ (4)
+where $I_{t}$ and $\hat{I}_{t}$ refer to the target image and the corresponding synthesized image, respectively.
+Note that $\hat{I}_{t}$ in the temporal-view photometric loss $\mathcal{L}_{temporal}$ is generated by projecting pixels from the source image using the coordinate index.
+In contrast, $\hat{I}_{t}$ in the cross-view photometric loss $\mathcal{L}_{cross}$ is derived from our proposed cross-view Gaussian splatting method.
+$\mathcal{L}_{semantic}$ is 2D semantic loss.
+The $\beta$ is set to 0.15 to balance the two loss terms.
+For $\mathcal{L}_{temporal}$ and $\mathcal{L}_{semantic}$, we recommend [54] for further details.
 Figure 3: Overlap mask in nuScenes [3] and DDAD [13].
+Method GT Occ. GT Pose mIoU* mIoU
+MonoScene [4] $\times$ 6.33 6.06
+BEVDet  [20] $\times$ 20.03 19.38
+BEVFormer [29] $\times$ 24.64 23.67
+OccFormer [56] $\times$ 22.39 21.93
+RenderOcc [37] $\times$ 24.53 23.93
+TPVFormer [22] $\times$ 28.69 27.83
+CTF-Occ [41] $\times$ 29.54 28.53
+SimpleOcc [11] $\times$ 7.99 7.05
+SelfOcc [23] $\times$ 10.54 9.30
+OccNeRF [54] $\times$ 10.81 9.54
+GaussianOcc $\times$ $\times$ 11.26 9.94
+Table 1: 3D occupancy prediction performance on the Occ3D-nuScenes dataset in mIoU metric.
+Since ‘other’ and ‘other flat’ classes are the invalid prompts for open-vocabulary models, we also calculate ‘mIoU*’ as the result ignoring the classes that do not consider these two classes during evaluation, while ‘mIoU’ is the original result.
+GT Occ.
+means using the ground truth occupancy label for the supervision.
+GT Pose is the ground truth pose from the sensor for self-supervised geometry learning.
+The full table is placed in the supplementary material with each category result.
+Method GT pose Occ. Abs Rel Sq Rel RMSE RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
+nuScenes [3]
+FSM [15] $\times$ $\times$ 0.297 - - - - - -
+FSM* [15] $\times$ $\times$ 0.319 7.534 7.860 0.362 0.716 0.874 0.931
+SurroundDepth [44] $\times$ $\times$ 0.280 4.401 7.467 0.364 0.661 0.844 0.917
+SA-FSM [50] $\times$ $\times$ 0.272 4.706 7.391 0.355 0.689 0.868 0.929
+VFF [27] $\times$ $\times$ 0.289 5.718 7.551 0.348 0.709 0.876 0.932
+R3D3 [38] $\times$ $\times$ 0.253 4.759 7.150 - 0.729 - -
+GaussianOcc $\ddagger$ $\times$ $\times$ 0.258 5.733 7.222 0.343 0.753 0.888 0.934
+SimpleOcc [11] 0.224 3.383 7.165 0.333 0.753 0.877 0.930
+OccNeRF [54] 0.202 2.883 6.697 0.319 0.768 0.882 0.931
+OccNeRF  [54] $\dagger$ 0.456 12.682 9.194 0.399 0.704 0.833 0.890
+SelfOcc [23] 0.215 2.743 6.706 0.316 0.753 0.875 0.932
+GaussianOcc $\times$ 0.211 3.115 7.131 0.326 0.762 0.878 0.931
+GaussianOcc $\dagger$ $\times$ 0.197 1.846 6.733 0.312 0.746 0.873 0.931
+DDAD [13]
+FSM*  [15] $\times$ $\times$ 0.228 4.409 13.433 0.342 0.687 0.870 0.932
+VFF [27] $\times$ $\times$ 0.218 3.660 13.327 0.339 0.674 0.862 0.932
+SurroundDepth [44] $\times$ $\times$ 0.208 3.371 12.977 0.330 0.693 0.871 0.934
+SA-FSM [50] $\times$ $\times$ 0.187 3.093 12.578 0.311 0.731 0.891 0.945
+R3D3 [38] $\times$ $\times$ 0.162 3.019 11.408 - 0.811 - -
+GaussianOcc $\ddagger$ $\times$ $\times$ 0.212 3.556 12.564 0.320 0.701 0.888 0.944
+GaussianOcc $\times$ 0.228 3.854 14.326 0.357 0.660 0.853 0.922
+Table 2: Comparisons for self-supervised multi-camera depth estimation on the nuScenes  [3] and DDAD datasets  [13].
+The results are averaged over all views without median scaling at test time.
+‘FSM*’ is the reproduced result in [27].
+GaussianOcc $\ddagger$ represents the depth estimation result from Stage 1.
+GaussianOcc $\dagger$ and OccNeRF $\dagger$ means the model trained with the semantic information.
+Occ.
+represents the ability of the method to predict the 3D occupancy.
+Figure 4: Visualization of the render depth map and 3D occupancy prediction on the nuScenes and DDAD datasets.
+4 Experiment
+4.1 Tasks, datasets, and metric
+nuScenes [3]: For 3D occupancy estimation, we utilize annotations from Occ3D-nuScenes [41], which includes 600 scenes for training and 150 scenes for evaluation.
+Each sample has a perception range of [-40m, -40m, -1m, 40m, 40m, 5.4m] with a voxel size of 0.4m.
+For fair comparison, we use the 2D semantic map provided by OccNeRF [54].
+We measure 3D occupancy estimation performance using the mean Intersection over Union (mIoU) metric.
+For depth estimation, we set the perception range in [-80m, -80m, -1m, 80m, 80m, 6m], while we clamp the ground truth to a range of 0.1m to 80m for evaluation, consistent with OccNeRF and SurroundDepth [44].
+We evaluate depth maps using error metrics (Abs Rel, Sq Rel, RMSE, RMSE log) and threshold accuracy metrics ($\delta$).
+DDAD [13]: Due to the lack of ground truth labels, we only present qualitative results for the fully self-supervised 3D occupancy estimation setting.
+Following the procedure on the nuScenes dataset, we obtain the 2D semantic labels according to the OccNeRF pipeline.
+For depth estimation, we also set the perception range in [-80m, -80m, -1m, 80m, 80m, 6m], but clamp the ground truth to a range of 0.1m to 200m for evaluation, consistent with SurroundDepth [44].
+Scale-aware training in [44, 27] Scale-aware training by ours RMSE
+Loss in [44] Loss in [27] GS loss Mask Erode Refine Abs Rel Sq Rel RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
+✓* 0.280 4.401 7.467 0.364 0.661 0.844 0.917
+0.672 25.405 11.999 0.568 0.419 0.808 0.878
+✓* 0.289 5.718 7.551 0.348 0.709 0.876 0.932
+0.285 6.046 7.514 0.342 0.702 0.865 0.931
+0.798 11.571 15.251 1.472 0.006 0.015 0.028
+0.293 7.127 7.536 0.376 0.743 0.876 0.923
+0.281 6.986 7.347 0.354 0.766 0.885 0.929
+0.258 5.733 7.222 0.343 0.753 0.888 0.934
+Table 3: Ablation study for scale-aware depth estimation on the nuScenes dataset [14].
+✓* means the result from the original paper and ✓means the result using New-CRFs [52] as the depth network the same as ours.
+GS loss means using the spatial context constraint by our proposed Gaussian splatting for projection.
+Mask represents using the mask-out strategy before the unprojection.
+Erode means the erode process to the binary overlap mask and Refine is the refinement of depth estimation network with 2 epochs by fixing the 6D pose network.
+4.2 Implementation details
+Network details: For U-Net architecture, we adapt New-CRFs [52] to predict the Gaussian attributes, which is based on the Swin Transformer [33].
+The 6D pose network is the same as that used in SurroundDepth [44].
+For the 2D-to-3D lifting, we follow the approach used in SimpleOcc [11].
+In the depth estimation benchmark, we use the network proposed by SimpleOcc, where the final output size is 256×256×16.
+In our Gaussian splatting setting, we further upsample the final output to 512×512×32 for improved performance since we observe that a finer voxel grid leads to a finer rendered depth map, which requires ignorable computational cost.
+For occupancy estimation, we use the same network as OccNeRF [54] to ensure a fair comparison.
+Training details: We propose a two-stage training process for fully self-supervised 3D occupancy estimation.
+In stage 1, we jointly train the depth estimation network and the 6D pose network, where we train the models for 8 epochs on the nuScenes dataset and 12 epochs on the DDAD dataset.
+In stage 2, we train the 3D occupancy network in a self-supervised manner with the 6D pose predicted from stage 1 rather than the ground truth pose used in OccNeRF [54].
+We train the models for 12 epochs on both the nuScenes and DDAD datasets.
+The optimizer and learning rate adjustment strategy follow those used in SimpleOcc [11] and OccNeRF [54].
+More implementation details (including code) are provided in the supplementary material.
+4.3 Main results
+3D occupancy estimation in nuScenes: As shown in Table 1, the proposed GaussianOcc achieves the best performance compared to other self-supervised methods.
+In particular, we do not require ground truth occupancy labels and ground truth 6D poses for training, thanks to the predicted 6D pose from the stage 1.
+In addition, we also achieve the good result on the RayIoU metric [31], with further details provided in the supplementary material.
+3D occupancy estimation in DDAD: Thanks to our fully self-supervised learning setting, we are able to perform 3D occupancy estimation on the DDAD dataset [13].
+To the best of our knowledge, our work is the first to introduce 3D occupancy estimation on this dataset, although we cannot provide quantitative results due to the lack of 3D occupancy label.
+We present visualization results in Figure 4.
+More results can be found in the supplementary material.
+Depth estimation: We present a comparison of depth estimation results in Table 2 for both the nuScenes and DDAD datasets.
+In stage 1, GaussianOcc $\ddagger$ achieves top performance on the nuScenes dataset and delivers competitive results on the DDAD dataset.
+It is important to note that methods such as SurroundDepth [44] and SA-FSM [50] rely on third-party module for the sparse depth supervision.
+Additionally, SA-FSM is preprint work and has not released the code.
+R3D3 [38] is a temporal offline refinement method that requires multi-frame optimization, as discussed in [23].
+In stage 2, which involves depth estimation from rendering, our method also achieves competitive results compared to those trained with ground truth poses.
+Besides, we observe that the rendered depth in stage two outperforms the depth results from stage 1 on the nuScenes dataset, whereas the opposite is true for the DDAD dataset.
+This discrepancy might be attributed to differences in perception range—80 meters in nuScenes versus 200 meters in DDAD.
+Moreover, an interesting phenomenon is that the semantic information is helpful for the depth estimation as indicated in GaussianOcc $\dagger$ whereas it worsens the result in OccNeRF $\dagger$.
+This shows the difference between the Gaussian splatting rendering and volume rendering.
+4.4 Ablation Study
+Scale-aware training: To demonstrate the effectiveness of the proposed scale-aware training using Gaussian Splatting, we conducted experiments comparing our method with existing approaches [44, 27] on Table 3.
+For a more fair comparison, we also implemented scale-aware training on [44, 27] that uses the same depth estimation network as ours [52], noted as $\checkmark$.
+We could reproduce the result on [27], but it is not good to training network with the sparse depth following [44].
+Our proposed scale-aware training method is better than [44, 27].
+Specifically, we observed that a naive implementation of Gaussian splatting without a masking strategy for cross-view rendering is ineffective because it would lead to a sub-optimal solution that the rendered image is still from the current view.
+Besides, to enhance performance, we introduced an erosion operation on the binary mask to purify it, excluding regions that may fall outside the overlap area.
+This step ensures better alignment during training.
+Finally, we refined the depth estimation by fixing the 6D pose network and disabling the cross-view loss.
+This refinement helps reduce artifacts at the edges of the overlap region.
+We present the cross-view splatting render result in Figure 5, where the rendering result in overlap regions is reasonable in the correct depth estimation from cross view that verifies our design.
+6D pose learning and training strategy: Considering the GT pose from the sensor is imperfect as a lack of vertical movement [36].
+We further evaluate the learned 6D pose quality by the self-supervised depth result at stage 2, where the experiments on Table 4 use the same network but with different pose constraints.
+Our result is competitive compared with GT pose, indicating that the predicted pose is of high quality.
+Furthermore, one stage training was ineffective in using the 6D pose from the jointly trained with 6D pose network because the cross-view loss in the 3D voxel space led to local optimization, which failed to generalize predictions to non-overlapping regions., as shown in Figure 6.
+However, this issue did not arise in the depth maps generated by the 2D decoder in stage 1, which is why we need the two-stage training process.
+Besides, our result is also better than [27], which is consistent the depth result on Table 3.
+Volume rendering and Splatting rendering: In Table 5, we compare the performance of volume rendering (VR) and splatting rendering (SR) in the voxel space.
+For each 3D Gaussian, we set a uniform scale $s$ for all Gaussians, considering the well-arranged positions of vertices within the voxel grid.
+Since the scale $s$ is consistent across all three dimensions, we set the rotation $\mathbf{R}$ as the identity matrix.
+We first explored the optimal scale and found that a scale of 0.1 produced the best results.
+Notably, the performance of VR and SR was very close when using this scale.
+The render depth maps are shown in Figure 6.
+The depth map by the volume rendering is smoother thanks to dense sampling and the depth maps from splatting tendering have the graininess effect, especially at the small scale factor.
+Pose and render types ablation study in occupancy task: In Table 6, we present an ablation study to evaluate the impact of using ground truth versus learned poses in the occupancy task, with training conducted using different rendering methods.
+The results show that splatting rendering achieves superior performance in both occupancy metrics (mIoU*) and depth metrics.
+One observation is that depth estimation results with the semantic learning in volume rendering are significantly worse than those in Table 2 whereas the proposed splatting rendering method maintains consistent performance, which suggests the splatting rendering in voxel grid contributes better 3D geometry.
+Rendering efficiency analysis: We analyze rendering efficiency in Table 7.
+The results show that volume rendering consumes 5 times more rendering time compared to splatting rendering.
+As resolution increases, both volume rendering time and GPU consumption rise significantly.
+In contrast, splatting rendering shows no significant increase in computational cost with higher resolutions, highlighting its efficiency and scalability.
+Bonus of the fully self-supervised setting: The fully self-supervised setting of our method could be a general pretraining solution for supervised learning.
+After the self-supervised training on the DDAD and nuScenes datasets, we further finetune the model with the 3D occupancy label from Occ3D-nuScenes [41].
+As shown in Table 8, experiments with self-supervised pretraining outperform the baseline.
+In particular, we find that pretraining on nuScenes is better than the DDAD dataset, which may own to the domain gap factors, such as differences in the scenarios (RGB images) and sensor configurations (camera extrinsics).
+Pose type Abs Rel Sq Rel RMSE $\delta<1.25$
+GT pose 0.214 3.362 7.127 0.771
+One stage training 0.946 17.008 16.397 0.103
+[27] 0.235 3.592 7.295 0.750
+Ours 0.211 3.115 7.131 0.762
+Table 4: Comparison of pose type for stage 2 training on depth estimation task [14].
+One stage training directly uses the cross-view loss to the rendered depth map.
+Apart from the GT pose, we also experiment the learned pose from [27] for comparison.
+Render type Abs Rel Sq Rel RMSE $\delta<1.25$
+VR 0.215 3.508 7.113 0.775
+SR ($s$ = 0.05) 0.223 3.694 7.246 0.761
+SR ($s$ = 0.1) 0.217 3.504 7.152 0.770
+SR ($s$ = 0.15) 0.217 3.406 7.204 0.763
+Table 5: Comparison of the render result between the volume rendering (VR) [54] and splatting rendering (SR, Ours) on depth estimation task [14].
+The coefficient $s$ means the scale of the 3D Gaussian.
+We use GT pose for this set of the ablation study.
+Pose type mIoU* Abs Rel Sq Rel $\delta<1.25$
+GT pose (VR) 10.81 0.456 12.682 0.704
+GT pose (SR) 11.30 0.225 4.339 0.787
+Learned pose (VR) 11.19 0.506 15.577 0.684
+Learned pose (SR) 11.26 0.197 1.846 0.746
+Table 6: Comparison of ground truth pose (GT Pose) and our learned pose (Two stages) on 3D occupancy estimation task [14] in volume rendering (VR) and splatting rendering (SR).
+Render Render resolution and time (s) Training
+type 180 $\times$ 320 240 $\times$ 520 360 $\times$ 640 time (h)
+VR $\approx$ 0.85 $\approx$ 1.57 N/A $\approx$ 2.68
+SR $\approx$ 0.17 $\approx$ 0.17 $\approx$ 0.17 $\approx$ 1
+Table 7: Comparison of rendering efficiency between volume rendering (VR) [54] and splatting rendering (SR, Ours) on 3D occupancy estimation task [14].
+The render time is calculated from surround 6 images.
+”N/A” indicates out-of-memory errors running in NVIDIA A 100 (40 GB).
+Training time is averaged per epoch.
+Method Pretraining setting mIoU
+Baseline None 37.29
+Self-supervised pretraining DDAD 37.40
+nuScenes 38.45
+Table 8: The study on SimpleOcc [11] with fully self-supervised pretraining.
+The baseline is directly training the model with 3D occupancy labels.
+The self-supervised pretraining is conducted on DDAD and nuScenes and then finetuned the model with 3D occupancy label.
+The number with bold typeface means the best.
+Figure 5: Visualization of the depth map, splatting render image and depth map from cross view in stage 1.
+Figure 6: The comparison for the depth map in the different setting, corresponding to the training strategy in Table 4 and rendering type in Table 5.
+5 Conclusion
+In this paper, we introduce GaussianOcc, a fully self-supervised and efficient method for 3D occupancy estimation.
+Through the carefully designed cross-view splatting rendering, we can accurately learn the real scale in depth and the 6D pose, enabling effective self-supervised 3D occupancy learning.
+Additionally, the proposed Gaussian splatting in voxel grids outperforms volume rendering in 3D occupancy estimation while reducing computational cost.
+As a bonus, the fully self-supervised setting of the proposed method could serve as the general pretrain solution to enhance the 3D occupancy estimation performance.
+We validate our method on the nuScenes and DDAD datasets, demonstrating its strong generalization ability to surround view scenes.
+6 More implementation details
+The detailed parameter setting in Gaussian attributes estimation network [52].
+During the joint depth and 6D pose training in stage 1, we predict the 3D Gaussian parameters alongside the 2D depth map.
+Since the Gaussian parameters are well-arranged in the 2D image plane prior to unprojection, we maintain equal scaling across all three dimensions of each 3D Gaussian and constrain the maximum scale to 0.02.
+Given that the scale $s$ is uniform across all dimensions, we set the rotation matrix $\mathbf{R}$ to the identity matrix.
+Additionally, we assign an opacity value of 1 to each 3D Gaussian, ensuring that every 2D depth value corresponds to a valid point in 3D space.
+We do not predict the color defined by SH coefficients $c$, while we directly use the source RGB image as the color map the same as in [57].
+The detailed parameter setting in voxel grid splatting rendering.
+We chose a fixed scale for each grid vertex to ensure a well-arranged structure that accurately models the 3D space.
+If we use a learnable scale, it may lead to a situation where the scale is small but the opacity is large, which may not be captured in the rendered depth map and semantic map but could still affect the 3D occupancy result.
+Therefore, using a fixed scale is simple and sufficient for optimization, as demonstrated in the results presented in the main paper.
+Since the scale $s\in\mathbb{R}^{3}_{+}$ are identical for both three dimensions, we do not need to predict the rotation $r\in\mathbb{R}^{4}$ and set it with the identical matrix is sufficient.
+Similar to the OccNeRF [54], we render the 2D feature map for the semantic regression, while we leverage the 3D Gaussian splatting rendering and OccNeRF uses volume rendering.
+The detailed parameter setting in training.
+We follow the training setting as OccNeRF [54], the resolution of input images and rendered depth maps are set as 384$\times$640 and 180$\times$320 respectively.
+All experiments are conducted on 8 NVIDIA A100 (40 GB).
+The detailed training strategy of the self-supervised pretraining setting.
+We first do the self-supervised training with 12 epochs with the learned pose from stage 1 and the 2D pseudo semantic label, which does not require the 3D occupancy label.
+Then, we finetune the model with 12 epochs with the 3D occupancy label.
+We add the RayIoU metric [31] in Table 11 for a comprehensive comparison.
+The detailed definition of depth map metric. Following the depth estimation task [44], we report the depth map evaluation with the following metrics,
+$\begin{split}\text{Abs Rel:}\frac{1}{|M|}\sum_{d\in M}\left|\hat{d}-d^{*}\right|/d^{*},\\ \text{Sq Rel:}\frac{1}{|M|}\sum_{d\in M}\left\|\hat{d}-d^{*}\right\|^{2}/d^{*},\\ \text{ RMSE: }\sqrt{\frac{1}{|M|}\sum_{d\in M}||\hat{d}-d^{*}\|^{2}},\\ \text{ RMSE log: }\sqrt{\frac{1}{|M|}\sum_{d\in M}||\log\hat{d}-\log d^{*}\|^{2}},\\ \delta<t:\%\text{ of }d\text{ s.t.
+}\max\left(\frac{\hat{d}}{d^{*}},\frac{d^{*}}{\hat{d}}\right)=\delta<t,\end{split}$ (5)
+where $M$ is the valid pixel, $\hat{d}$ is the ground truth depth and $d^{*}$ is the predicted depth.
+7 More experiment results
+mIoU metric. Due to the limited space in the main paper, the full table of mIoU results is presented in Table 9 for reference.
+RayIoU metric.
+In addition to the mIoU metric for 3D occupancy estimation, we also evaluate our method with a novel metric, RayIoU, introduced by the recent work [31].
+The RayIoU is a ray-based evaluation metric that resolves the inconsistency penalty along the depth axis introduced in the traditional voxel-level mIoU criteria.
+As shown in Table 10, our approach also outperforms OccNeRF [54] in this metric as well.
+It’s important to note that the FPS is calculated excluding rendering time.
+Since GaussianOcc and OccNeRF utilize the same network architecture, they share the same inference time when the rendering process is not taken into account.
+More visualization. We provide more visualization for nuScenes dataset in Figure 7 and DDAD dataset in Figure 8. Please check the attached videos for sequence visualization.
+More analysis on different supervision types.
+In Figure 9, we present visualizations of different supervision types.
+These visualizations highlight key differences in the results for the invisible regions (marked with red rectangles) and the rendered depth quality (marked with green rectangles).
+Experiments (1) and (2) involve supervision using ground truth (GT) occupancy labels. Specifically:
+Experiment (1) is trained without the visible mask provided by Occ3D-nuScenes [41], which defines the visibility of the occupancy labels.
+Without this mask, the invisible regions are treated as empty, and the loss function is applied to these regions as well.
+Experiment (2), on the other hand, excludes the loss computation in invisible regions.
+From the results, we observe that in Experiment (1), the model tends to predict empty values for invisible regions due to empty loss penalty.
+In contrast, Experiment (2), by ignoring the loss in invisible areas, shows more non-empty predictions in these regions.
+Self-supervised experiments (3) and (4) rely on rendering techniques, which inherently cannot optimize predictions in invisible regions.
+This limitation leads to non-empty predictions in the red-highlighted areas.
+Notably, Experiment (4) frequently predicts invisible regions as related to foreground categories, as shown in the dark rectangles.
+Conversely, Experiment (3) demonstrates a consistent tendency to classify invisible regions as man-made structures, likely because the surrounding environment predominantly consists of man-made elements.
+Additionally, the depth maps in the green-highlighted areas show that our Gaussian splatting rendering produces higher-quality depth predictions compared to volume rendering, consistent with the quantitative results in Table 6 of the main paper.
+Since Experiments (1) and (2) do not involve rendering-based training, they fail to produce reasonable depth predictions.
+8 Limitation and future work
+The proposed method achieves reasonable predictions in most scenes; however, we observe that some cases still present challenges, as shown in Figure 10.
+Specifically, in the DDAD dataset, incorrect predictions occur in the back camera in certain situations as marked with the red circle, where the drivable surface is mistakenly projected into the car due to extensive self-occlusion.
+Notably, this issue is absent in the nuScenes dataset, which has less self-occlusion.
+We believe that this problem could be mitigated with better 2D semantic maps for supervision, which warrants further investigation.
+The proposed method is for the surround view setting which is not suitable for the monocular images and the two-stage training procedure is a little bit time-consuming.
+Therefore, in the future, we will further investigate to apply the proposed method for one-stage training.
+Additionally, in stage 1, we leverage the spatial cross-view constraint through the proposed Gaussian splatting method.
+In the future, we aim to explore its potential benefits for temporal view synthesis as well.
 Method
@@
 GT Pose
-mIoU mIoU*
+mIoU* mIoU
 $\blacksquare$ barrier
@@
 Table 9: 3D occupancy prediction performance on the Occ3D-nuScenes dataset in mIoU metric.
-Since ‘other’ and ‘other flat’ classes are the invalid prompts for open-vocabulary models, we also calculate ‘mIoU’ as the result ignoring the classes that do not consider these two classes during evaluation, while ‘mIoU*’ is the original result.
+Since ‘other’ and ‘other flat’ classes are the invalid prompts for open-vocabulary models, we also calculate ‘mIoU*’ as the result ignoring the classes that do not consider these two classes during evaluation, while ‘mIoU’ is the original result.
 GT Occ.
-means using the ground truth occupancy label for the supervision.
+refers to the use of the ground truth occupancy label for supervision.
 GT Pose is the ground truth pose from the sensor for self-supervised geometry learning.
@@
 GT Pose
-Backbone Input Size Epoch RayIoU RayIoU1m, 2m, 4m mIoU* FPS
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Backbone Input Size Epoch RayIoU RayIoU1m, 2m, 4m mIoU FPS
+BEVFormer (4f) [29] $\times$ R101 1600$\times$900 24 32.4 26.1 32.9 38.0 39.2 3.0
+RenderOcc [37] $\times$ Swin-B 1408$\times$512 12 19.5 13.4 19.6 25.5 24.4 -
+SimpleOcc [11] $\times$ R101 672$\times$336 12 28.2 22.3 28.7 33.7 37.3 9.7
+BEVDet-Occ (2f) [19] $\times$ R50 704$\times$256 90 29.6 23.6 30.0 35.1 36.1 2.6
+BEVDet-Occ-Long (8f) $\times$ R50 704$\times$384 90 32.6 26.6 33.1 38.2 39.3 0.8
+FB-Occ (16f) [30] $\times$ R50 704$\times$256 90 33.5 26.7 34.1 39.7 39.1 10.3
+SparseOcc (8f) $\times$ R50 704$\times$256 24 34.0 28.0 34.7 39.4 30.1 17.3
+SparseOcc (16f) $\times$ R50 704$\times$256 48 36.1 30.2 36.8 41.2 30.9 12.5
+OccNeRF [54] $\times$ R101 640$\times$384 12 10.49 6.93 10.28 14.26 9.54 10.8
+GaussianOcc $\times$ $\times$ R101 640$\times$384 12 11.85 8.69 11.90 14.95 9.94 10.8
 Table 10: 3D Occupancy prediction performance on the Occ3D-nuScenes dataset in RayIoU metric.
@@
 “8f” and “16f” mean fusing temporal information from 8 or 16 frames.
-mIoU* is the mean Intersection over Union for all categories.
+mIoU is the mean Intersection over Union for all categories.
 FPS means frame per second for each method, which is measured on a Tesla A100 GPU.
-Method GT pose Occ. Abs Rel Sq Rel RMSE RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
-nuScenes [3]
-FSM [16] $\times$ $\times$ 0.297 - - - - - -
-FSM* [16] $\times$ $\times$ 0.319 7.534 7.860 0.362 0.716 0.874 0.931
-SurroundDepth [44] $\times$ $\times$ 0.280 4.401 7.467 0.364 0.661 0.844 0.917
-SA-FSM [50] $\times$ $\times$ 0.272 4.706 7.391 0.355 0.689 0.868 0.929
-VFF [28] $\times$ $\times$ 0.289 5.718 7.551 0.348 0.709 0.876 0.932
-R3D3 [38] $\times$ $\times$ 0.253 4.759 7.150 - 0.729 - -
-GaussianOcc $\ddagger$ $\times$ $\times$ 0.258 5.733 7.222 0.343 0.753 0.888 0.934
-SimpleOcc [12] 0.224 3.383 7.165 0.333 0.753 0.877 0.930
-OccNeRF [54] 0.202 2.883 6.697 0.319 0.768 0.882 0.931
-SelfOcc [24] 0.215 2.743 6.706 0.316 0.753 0.875 0.932
-GaussianOcc $\times$ 0.211 3.115 7.131 0.326 0.762 0.878 0.931
-DDAD [14]
-FSM*  [16] $\times$ $\times$ 0.228 4.409 13.433 0.342 0.687 0.870 0.932
-VFF [28] $\times$ $\times$ 0.218 3.660 13.327 0.339 0.674 0.862 0.932
-SurroundDepth [44] $\times$ $\times$ 0.208 3.371 12.977 0.330 0.693 0.871 0.934
-SA-FSM [50] $\times$ $\times$ 0.187 3.093 12.578 0.311 0.731 0.891 0.945
-R3D3 [38] $\times$ $\times$ 0.162 3.019 11.408 - 0.811 - -
-GaussianOcc $\ddagger$ $\times$ $\times$ 0.212 3.556 12.564 0.320 0.701 0.888 0.944
-GaussianOcc $\times$ 0.228 3.854 14.326 0.357 0.660 0.853 0.922
-Table 3: Comparisons for self-supervised multi-camera depth estimation on the nuScenes  [3] and DDAD datasets  [14].
-The results are averaged over all views without median scaling at test time.
-‘FSM*’ is the reproduced result in [28].
-GaussianOcc $\ddagger$ represents the depth estimation result from Stage 1.
-Occ.
-represents the ability of the method to predict the 3D occupancy.
-Figure 4: The visualization of the render depth map and 3D occupancy prediction on nuScenes and DDAD dataset.
-4 Experiment
-4.1 Tasks, Datasets, and Metric
-nuScenes: We perform 3D occupancy estimation and depth estimation using the nuScenes dataset [3].
-For 3D occupancy estimation, we utilize annotations from Occ3D-nuScenes [41], which includes 600 scenes for training and 150 scenes for evaluation.
-Each sample has a perception range of [-40m, -40m, -1m, 40m, 40m, 5.4m] with a voxel size of 0.4m.
-For fair comparison, we use the 2D semantic map provided by OccNeRF [54].
-We measure 3D occupancy estimation performance using the mean Intersection over Union (mIoU) metric.
-For depth estimation, we set the perception range in [-80m, -80m, -1m, 80m, 80m, 6m], while we clamp the ground truth to a range of 0.1m to 80m for evaluation, consistent with OccNeRF and SurroundDepth [44].
-We evaluate depth maps using error metrics (Abs Rel, Sq Rel, RMSE, RMSE log) and threshold accuracy metrics ($\delta$).
-DDAD: We conduct 3D occupancy estimation and depth estimation using the DDAD dataset [14].
-Due to the lack of ground truth labels, we only present qualitative results for the fully self-supervised 3D occupancy estimation setting.
-Following the procedure used with the nuScenes dataset, we obtain the 2D semantic labels as per the OccNeRF pipeline.
-For depth estimation, we also set the perception range in [-80m, -80m, -1m, 80m, 80m, 6m], but clamp the ground truth to a range of 0.1m to 200m for evaluation, consistent with SurroundDepth [44].
-Scale aware training in [44] Scale aware training by ours RMSE
-Spatial loss SFM loss GS loss Mask Erode Refine Abs Rel Sq Rel RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
-✓* 0.970 13.702 17.931 3.654 0.000 0.000 0.001
-0.921 30.393 14.444 0.808 0.057 0.187 0.667
-✓* ✓* 0.280 4.401 7.467 0.364 0.661 0.844 0.917
-0.672 25.405 11.999 0.568 0.419 0.808 0.878
-0.798 11.571 15.251 1.472 0.006 0.015 0.028
-0.293 7.127 7.536 0.376 0.743 0.876 0.923
-0.281 6.986 7.347 0.354 0.766 0.885 0.929
-0.258 5.733 7.222 0.343 0.753 0.888 0.934
-Table 4: Ablation study for scale-aware depth estimation on the Nuscene dataset [15].
-The scores are averaged over all views without median-scaling at test time.
-✓* means the result from the original paper and ✓means the result from our implementation in New-CRFs network [52].
-Spatial loss denotes the spatial context loss used in [44] and SFM loss denotes the sparse depth pretrain from the off-line Structure-from-Motion process.
-GS loss means using the spatial context constraint by our proposed Gaussian splatting for projection.
-Mask represents using the mask-out strategy before the unprojection.
-Erode means use the erode process to the binary overlap mask and Refine is the refinement of the depth estimation network by fixing the 6D pose net.
-4.2 Implementation details
-Network details: For U-Net architecture, we adopt the New-CRFs [52] as the depth estimation network, which is based on the Swin Transformer [34].
-The 6D pose network is the same as that used in SurroundDepth [44].
-For the 2D to 3D lifting, we follow the approach used in SimpleOcc [12].
-In depth estimation benchmark, we use the network proposed by SimpleOcc, where the final output size is 256×256×16.
-In our Gaussian splatting setting, we further upsample the final output to 512×512×32 for improved performance, which only requires minimal additional computational cost.
-For occupancy estimation, we use the same network as OccNeRF [54] to ensure a fair comparison.
-Training details: We propose a two-stage training process for fully self-supervised 3D occupancy estimation.
-In the first stage, we jointly train the depth estimation network and the 6D pose network.
-In the second stage, we train the 3D occupancy network without optimizing the 6D pose network.
-During the first stage, we train the models for 8 epochs on the nuScenes dataset and 12 epochs on the DDAD dataset.
-In the second stage, we train the models for 12 epochs on both the nuScenes and DDAD datasets.
-The optimizer and learning rate adjustment strategy are the same as those used in SimpleOcc [12] and OccNeRF [54].
-4.3 Main results
-3D occupancy estimation in mIoU metric: As shown in Table 1, the proposed GaussianOcc achieves the best performance compared with other self-supervised methods, without the need for ground truth occupancy labels and ground truth 6D poses for training.
-It’s worth noting that SelfOcc [24] differs slightly from methods like SimpleOcc, OccNeRF, and our approach.
-Unlike these methods, SelfOcc does not predict semantic information directly in the 3D occupancy space.
-Instead, it attaches 2D semantic information to the 3D occupied voxels, where the 2D semantic data is sourced from a third-party open-vocabulary model [55].
-3D occupancy estimation in RayIoU metric: We also compare our method using a newly proposed metric, RayIoU, introduced by the recent work [32].
-The RayIoU is a ray-based evaluation metric to solve the inconsistency penalty along the depth axis raised in traditional voxel-level mIoU criteria.
-As shown in Table 2, our approach also outperforms OccNeRF [54].
-It’s important to note that the FPS is calculated excluding rendering time.
-Since GaussianOcc and OccNeRF utilize the same network architecture, they share the same inference time when the rendering process is not taken into account.
-3D occupancy estimation in DDAD: Thanks to our fully self-supervised learning approach, we are able to perform 3D occupancy estimation on a similar dataset, DDAD [14].
-To the best of our knowledge, our work is the first to introduce 3D occupancy estimation on this dataset, although we cannot provide quantitative results due to the lack of 3D annotations.
-We present the visualization results in Figure 4 and more results could be found in the supplement material.
-Depth estimation: We present a comparison of depth estimation results in Table 3 for both the nuScenes and DDAD datasets.
-In the first stage of monocular depth estimation, GaussianOcc $\ddagger$ achieves top performance on the nuScenes dataset and delivers competitive results on the DDAD dataset.
-It is important to note that methods such as SurroundDepth [44] and SA-FSM [50] rely on third-party sparse point cloud extraction.
-Additionally, R3D3 [38] is a temporal offline refinement method that requires multi-frame optimization, as discussed in [24].
-In the second stage, which involves depth estimation from rendering, our method also achieves competitive results compared to those trained with ground truth poses.
-Besides, we observe that the rendered depth in stage two outperforms the depth results from stage 1 on the nuScenes dataset, whereas the opposite is true for the DDAD dataset.
-This discrepancy might be attributed to differences in perception range—80 meters in nuScenes versus 200 meters in DDAD.
-4.4 Ablation Study
-Scale aware training: To demonstrate the effectiveness of the proposed scale-aware training using Gaussian Splatting, we conducted experiments comparing our method with existing approaches that use SFM loss for training [44].
-As indicated in Table 4, we implemented scale-aware training from [44] on our network [52], but the results were suboptimal, and our proposed method outperformed the results reported in the original paper.
-In our method, we observed that a naive implementation of Gaussian splatting without a masking strategy is ineffective because it lacks cross-view rendering, which is essential for applying cross-view constraints.
-To enhance performance, we introduced an erosion operation on the binary mask to purify it, excluding regions that may fall outside the overlap area.
-This step ensures better alignment during training.
-Finally, we refined the depth estimation by fixing the 6D pose network and disabling the cross-view loss.
-This refinement helps reduce artifacts at the edges of the overlap region.
-We present the cross-view splatting render result in Figure 5, where the rendering result in overlap regions is reasonable in the correct depth estimation from cross view.
-6D pose and training strategy: Building on the success of scale-aware training, we compared the depth estimation performance of our proposed two-stage training method with that of models trained using ground truth (GT) pose.
-As shown in Table 5, our two-stage approach delivers competitive results when compared to training with GT pose, indicating that the predicted pose is of high quality.
-Additionally, the GT pose provided by the nuScenes dataset is limited by sensor inaccuracies, particularly a lack of vertical movement.
-In contrast, the learned 6D pose in our method does not suffer from this limitation.
-Furthermore, in our one-stage experiment, we found that this approach was not effective.
-The cross-view loss in the 3D voxel space led to local optimization, which failed to generalize predictions to non-overlapping regions as indicated in Figure 6.
-However, this issue did not arise in the depth maps generated by the 2D decoder in stage one.
-Volume rendering and Splatting rendering: In Table 6, we examine the performance of volume rendering (VR) and splatting rendering (SR) in the voxel space.
-For the scale of each 3D Gaussian, we set a uniform scale $s$ for all Gaussians, considering the well-arranged positions of vertices within the voxel grid.
-Since the scale $s$ is consistent across all three dimensions, we set the rotation $\mathbf{R}$ as the identity matrix.
-We first explored the optimal scale and found that a scale of 0.1 produced the best results.
-Notably, the performance of VR and SR was very close when using this scale.
-The render depth maps are shown in Figure 6.
-The depth map by the volume rendering is smoother thanks to dense sampling and the depth maps with splatting tendering have the graininess effect, especially at the small scale factor.
-Pose and render types ablation study in occupancy task: In Table 7, we conduct an ablation study to evaluate the impact of using ground truth (GT) versus learned poses in the occupancy task, with training conducted using different rendering methods.
-The results show that splatting rendering achieves superior performance in both occupancy metrics (mIoU) and depth metrics.
-One observation is that depth estimation results from 3D semantic learning with volume rendering are significantly worse than those in Table 3 and the proposed splatting rendering method maintains consistent performance, which indicates the splatting rendering in voxel grid may contribute better 3D geometry.
-Rendering efficiency analysis: Finally, we analyze rendering efficiency in Table 8.
-The results indicate that volume rendering consumes 5 times more rendering time compared to splatting rendering, and as the resolution increases, both rendering time and GPU consumption rise significantly.
-In contrast, splatting rendering shows no significant increase in computational cost with higher resolutions, highlighting its efficiency and scalability.
-Pose type Abs Rel Sq Rel RMSE $\delta<1.25$
-GT pose 0.214 3.362 7.127 0.771
-Two stages 0.211 3.115 7.131 0.762
-One stage 0.946 17.008 16.397 0.103
-Table 5: Comparison of the ground truth pose (GT pose) and our learned pose (Two stages) on the depth estimation task [15].
-We also conducted the experiment that directly uses the cross-view loss to render a depth map (One stage).
-Render type Abs Rel Sq Rel RMSE $\delta<1.25$
-VR 0.215 3.508 7.113 0.775
-SR ($s$ = 0.05) 0.223 3.694 7.246 0.761
-SR ($s$ = 0.1) 0.217 3.504 7.152 0.770
-SR ($s$ = 0.15) 0.217 3.406 7.204 0.763
-Table 6: Comparison of the render result between the volume rendering (VR) [54] and splatting rendering (SR, Ours) on depth estimation task [15].
-The coefficient $s$ means the scale of the 3D Gaussian.
-We use GT pose for this set of the ablation study.
-Pose type mIoU Abs Rel Sq Rel $\delta<1.25$
-GT pose (VR) 10.81 0.456 12.682 0.704
-GT pose (SR) 11.30 0.225 4.339 0.787
-Learned pose (VR) 11.19 0.506 15.577 0.684
-Learned pose (SR) 11.43 0.219 3.625 0.778
-Table 7: Comparison of ground truth pose (GT Pose) and our learned pose (Two stages) on 3D occupancy estimation task [15] in volume rendering (VR) and splatting rendering (SR).
-Render Render resolution and time (s) Training
-type 180 $\times$ 320 240 $\times$ 520 360 $\times$ 640 time (h)
-VR $\approx$ 0.85 $\approx$ 1.57 N/A $\approx$ 2.68
-SR $\approx$ 0.17 $\approx$ 0.17 $\approx$ 0.17 $\approx$ 1
-Table 8: Comparison of rendering efficiency between volume rendering (VR) [54] and splatting rendering (SR, Ours) on 3D occupancy estimation task [15].
-The render time is calculated from surround 6 images.
-”N/A” indicates out-of-memory errors running in NVIDIA A 100 (40 GB).
-Training time is averaged per epoch.
-Figure 5: The visualization for the depth map, splatting render image and depth map from cross view in stage 1.
-Figure 6: The comparison for the depth map in the different setting, corresponding to the training strategy in Table 5 and rendering type in Table 6.
-5 Conclusion
-In this paper, we introduce GaussianOcc, a fully self-supervised and efficient approach for 3D occupancy estimation.
-Through the carefully designed cross-view splatting rendering, we can accurately learn the real scale in depth and the 6D pose net, enabling effective self-supervised 3D occupancy learning.
-Additionally, the proposed Gaussian splatting in voxel grids outperforms volume rendering in 3D occupancy estimation while reducing computational cost.
-We validate our method on the nuScenes and DDAD datasets, demonstrating its strong generalization ability to surround view scenes.
-Appendix
-Appendix A More implementation details
-The detailed parameter setting in depth estimation network [52].
-During the joint depth and 6D pose training in stage 1, we predict the 3D Gaussian parameters alongside the 2D depth map.
-Since the Gaussian parameters are well-arranged in the 2D image plane prior to unprojection, we maintain equal scaling across all three dimensions of each 3D Gaussian and constrain the maximum scale to 0.02.
-Given that the scale $s$ is uniform across all dimensions, we set the rotation matrix $\mathbf{R}$ to the identity matrix.
-Additionally, we assign an opacity value of 1 to each 3D Gaussian, ensuring that every 2D depth value corresponds to a valid point in 3D space.
-The detailed parameter setting in voxel grid splatting rendering.
-We chose a fixed scale for each grid vertex to ensure a well-arranged structure that accurately models the 3D space.
-If we were to use a learnable scale, it could lead to a situation where the scale is small but the opacity is large, which might not be captured in the rendered depth map and semantic map but could still impact the 3D occupancy result.
-Therefore, using a fixed scale is sufficient for optimization, as demonstrated in the results presented in the main paper.
-[…]
-Detailed definition of depth map metric. Following the depth estimation task [44], we report the depth map evaluation with the following metrics,
-[…]
-where $M$ is the valid pixel, $\hat{d}$ is the ground truth depth and $d^{*}$ is the predicted depth.
-Appendix B More experiment results
-More visualization. We provide more visualization for nuScenes dataset in Figure A and DDAD dataset in Figure B. Please check the attached videos for sequence visualization.
-Appendix C Limitation and future work
-The proposed method achieves reasonable predictions in most scenes; however, we observe that some cases still present challenges, as shown in Figure C.
-Specifically, in the DDAD dataset, incorrect predictions occur in the back camera in certain situations as marked with the red circle, where the drivable surface is mistakenly projected into the car due to extensive self-occlusion.
-Notably, this issue is absent in the nuScenes dataset, which has less self-occlusion.
-We believe that this problem could be mitigated with better 2D semantic maps for supervision, which warrants further investigation.
-Additionally, we plan to explore the use of splatting rendering for 3D occupancy flow in future work.
-Figure A: The visualization of the render depth map and 3D occupancy prediction on nuScenes dataset.
-Figure B: The visualization of the render depth map and 3D occupancy prediction on DDAD dataset
-Figure C: Some wrong predictions due to the large self-occlusion on DDAD dataset.
+Figure 7: The visualization of the render depth map and 3D occupancy prediction on nuScenes dataset.
+Figure 8: The visualization of the render depth map and 3D occupancy prediction on DDAD dataset
+Figure 9: The visualization of the different supervision types (1-4) comparison on nuScenes dataset.
+Method Pretraining setting mIoU RayIoU RayIoU1m, 2m, 4m
+Baseline None 37.29 28.2 22.3 28.7 33.7
+Self-supervised pretrain DDAD 37.40 28.7 22.9 29.1 34.0
+nuScenes 38.45 29.9 23.9 30.4 35.5
+Table 11: The study on SimpleOcc [11] with fully self-supervised pretrain.
+The baseline is directly training the model with 3D occupancy label.
+The self-supervised pretraining is conducted on DDAD and nuScenes and then finetuned the model with 3D occupancy label.
+The number with bold typeface means the best.
+Figure 10: Some wrong predictions due to the large self-occlusion on DDAD dataset.

### v4

@@
 Abstract
-We introduce GaussianOcc, a systematic method that investigates two usages of Gaussian splatting for fully self-supervised and efficient 3D occupancy estimation in surround views.
-First, traditional methods for self-supervised 3D occupancy estimation still require ground truth 6D poses from sensors during training.
+We introduce GaussianOcc, a systematic method that investigates Gaussian splatting for fully self-supervised and efficient 3D occupancy estimation in surround views.
+First, traditional methods for self-supervised 3D occupancy estimation still require ground truth 6D ego pose from sensors during training.
 To address this limitation, we propose Gaussian Splatting for Projection (GSP) module to provide accurate scale information for fully self-supervised training from adjacent view projection.
@@
 We propose Gaussian Splatting from Voxel space (GSV) to leverage the fast rendering properties of Gaussian splatting.
-As a result, the proposed GaussianOcc method enables fully self-supervised (no ground-truth pose) 3D occupancy estimation in competitive performance with low computational cost (2.7 times faster in training and 5 times faster in rendering).
+As a result, the proposed GaussianOcc method enables fully self-supervised (no ground truth ego pose) 3D occupancy estimation in competitive performance with low computational cost (2.7 times faster in training and 5 times faster in rendering).
 The relevant code is available in https://github.com/GANWANSHUI/GaussianOcc.git.
 Abstract
-In this supplementary material, we provide more implementation details, experiment results, and further discussion on the limitations and future work.
+In this supplementary material, we provide more implementation details, experiment results with analysis, and further discussion on the limitations and future work.
 1 Introduction
 Figure 1: Problem setting of GaussianOcc.
-Given a surround image sequence and its corresponding 2D semantic annotation, GaussianOcc is able to perform 3D occupancy estimation without the need for ground truth occupancy label and ground truth 6D pose for training.
-In recent years, both vision-centric and grid-centric perception approaches have gained significant attention in the autonomous driving industry and academia [34, 39].
-Among these, surround view 3D occupancy estimation [31, 45, 51, 48] has emerged as a core perception task and a promising alternative to bird’s-eye view (BEV) methods [9, 28, 29].
+Given a surround image sequence, the spatial camera extrinsic and its corresponding 2D semantic annotation, GaussianOcc is able to perform 3D occupancy estimation without the need for ground truth occupancy label and ground truth 6D ego pose for training.
+Surround view 3D occupancy estimation [31, 45, 51, 48, 34, 39, 56] has emerged as a core perception task and a promising alternative to bird’s-eye view (BEV) methods [9, 28, 29].
 To facilitate 3D occupancy estimation, several benchmarks have been developed for supervised training [45, 43, 41, 42], though these require substantial effort in 3D annotation.
-To reduce the burden of annotation, self-supervised [11, 54, 23, 5, 17] and weakly-supervised [37] learning approaches based on volume rendering have been proposed [35, 46, 10].
-Volume rendering allows 3D representation learning using 2D supervision signals, such as 2D semantic maps and depth maps, thereby eliminating the need for extensive 3D annotation during training.
-Existing methods [54, 23] achieve self-supervised learning through volume rendering, where the 2D semantic map supervision is derived from open-vocabulary semantic segmentation [55], and the depth map supervision is obtained from self-supervised depth estimation [12].
+To reduce the burden of 3D annotation, self-supervised [11, 54, 22, 5, 16] and weakly-supervised [37] learning approaches based on volume rendering have been proposed [35, 46, 10].
+Volume rendering allows 3D representation learning using 2D supervision signals, such as 2D semantic maps and depth maps, thereby eliminating the need for extensive 3D annotation.
+Existing methods [54, 22] achieve self-supervised learning through volume rendering, where the 2D semantic map supervision is derived from open-vocabulary semantic segmentation [55], and the depth map supervision is obtained from self-supervised depth estimation [11, 12].
 However, these approaches face two significant limitations.
-First, volume rendering is performed at real-world scale, which requires the availability of ground truth 6D pose to calculate the multi-view photometric loss across sequential images.
+First, volume rendering is performed at real-world scale, which requires the ground truth 6D ego pose to calculate the multi-view photometric loss across sequential images.
 Second, the inefficiency in volume rendering poses a challenge, the same as in novel view synthesis tasks [26, 1, 17], due to the dense sampling operation required.
 These limitations impede the development of a more general and efficient paradigm for self-supervised 3D occupancy estimation.
-To address the aforementioned limitations, we propose a fully self-supervised and efficient approach to 3D occupancy estimation using Gaussian splatting [26, 1].
+To address the aforementioned limitations, we propose a fully self-supervised and efficient approach to 3D occupancy estimation with dedicated designs based on Gaussian splatting [26, 1].
 Specifically, we introduce the use of Gaussian splatting to perform cross-view splatting, where the rendered image constructs a cross-view loss that provides scale information during joint training with the 6D pose network.
-This eliminates the need for ground truth 6D pose during training.
+This eliminates the need for ground truth 6D ego pose during training.
 To improve rendering efficiency, we move away from the dense sampling required in traditional volume rendering.
 Instead, we propose performing Gaussian splatting directly from the 3D voxel space.
-In this approach, each vertex in the voxel grid is treated as a 3D Gaussian, and we optimize the attributes of these Gaussians—such as semantic features and opacity—directly within the voxel space.
-Through this novel approach, our proposed method, GaussianOcc, makes a progress toward fully self-supervised and efficient 3D occupancy estimation, as outlined in Figure 1.
+In this approach, each vertex in the voxel grid is treated as a 3D Gaussian, and we optimize the attributes of these Gaussians—such as semantic and opacity—directly within the voxel space.
+Through the above innovative approach, our proposed method makes progress toward fully self-supervised and efficient 3D occupancy estimation, as outlined in Figure 1.
 In summary, our core contributions are as follows:
   • We introduce the first fully self-supervised method for efficient surrounding-view 3D occupancy estimation, featuring the exploration of Gaussian splatting.
-  • We propose Gaussian splatting for cross-view projection module, which can provide scale information to get rid of the need of ground truth 6D pose during the training.
+  • We propose Gaussian splatting for cross-view projection module, which can provide scale information to get rid of the need of ground truth 6D ego pose during the training.
   • We propose Gaussian splatting from voxel space module, achieving competitive performance with 2.7 times faster training and 5 times faster rendering compared to the previous works with volume rendering.
@@
 In Stage 1, we train a scale-aware 6D pose network, using a U-Net architecture to predict Gaussian attributes in the 2D image grid space for cross-view Gaussian splatting.
-This approach provides scale information in the joint training with the 6D pose network.
+This approach provides scale information in the joint training with the 6D pose net.
 Based on the 6D pose from Stage 1, we perform self-supervised 3D occupancy estimation in Stage 2, where we lift the 2D features to a 3D voxel space and propose voxel grid Gaussian splatting for fast rendering.
@@
 In addition, we investigate the Gaussian splatting on the voxel space for faster rendering compared with the previous volume rendering-based works [11, 54, 22].
-Note that, two recent works, GaussianFormer [24] and GaussianBeV [6], are closely related to ours in their focus on 3D occupancy estimation and BEV prediction.
+Note that, recent works, GaussianFormer [23], GaussTR [24] and GaussianBeV [6], are related to ours in their focus on 3D occupancy estimation and BEV prediction.
 However, our exploration diverges by focusing on two new properties that Gaussian splatting can contribute to occupancy estimation: scale-aware training and faster rendering.
@@
 The original Gaussian splatting is for scene-specific, fast 3D novel view synthesis, where the attributes of Gaussian points are optimized by the multi-view constraint.
-In contrast, in this work, we explore a setting where Gaussian attributes are well-aligned in both 2D and 3D grids, allowing us to benefit from the Gaussian splatting rendering for the scale-aware training by cross-view constraint and faster rendering on voxel grids as illustrated in Figure 2.
-3.2 Scale-ware training by Gaussian Splatting
+Differently, we study fully self-supervised and efficient 3D occupancy estimation by exploring the Gaussian attributes that are well-aligned in both 2D and 3D grids.
+Our design allows us to benefit from the Gaussian splatting rendering for the scale-aware training by cross-view constraint and faster rendering on voxel grids, as illustrated in Figure 2.
+3.2 Scale-aware training by Gaussian Splatting
 Scale from spatial camera rig: Similar to the previous work [14, 44, 27], the scale information is from the surround camera rig.
-Specifically, the real-world scale can be obtained by leveraging camera extrinsic matrices which is to use spatial photometric loss in the overlap region between two adjacent views, i.e., warping ${I}^{i}_{t}$ to ${I}^{j}_{t}$:
-$p^{i\rightarrow j}_{t}=K^{j}(T^{j})^{-1}T^{i}D_{t}^{i}(K^{i})^{-1}p^{i}_{t}$ (1)
+Specifically, the real-world scale can be obtained by leveraging camera extrinsic matrices, which is to use spatial photometric loss in the overlap region between two adjacent views, i.e., warping ${I}^{i}_{t}$ to ${I}^{j}_{t}$:
+$p^{i\rightarrow j}_{t}=K^{j}(T^{j})^{-1}T^{i}D_{t}^{i}(K^{i})^{-1}p^{i}_{t},$ (1)
 where $K^{i},T^{i}$ are the intrinsic and extrinsic matrices of $i$-th camera, $D_{t}^{i}$ is the predicted depth map of $i$-th camera, $p_{t}$ is the corresponding pixel during the warping.
-The warping operation is achieved by bilinear interpolation with the corresponding $p^{i\rightarrow j}_{t}$.
-However, as pointed out by [44], the mapping $p^{i\rightarrow j}_{t}$ can easily go out of image bounds with the weak supervision signal in such a small overlap region.
-Therefore, apart from the spatial loss, [44] proposes to facilitate the Structure-from-Motion (SFM) to extract sparse depth information for the direct depth supervision to provide a stronger supervision signal, but it is time-consuming and not straightforward.
+The warping operation is achieved by direct bilinear interpolation with the corresponding $p^{i\rightarrow j}_{t}$.
+However, as pointed out by [44], the mapping in such a small overlap region, $p^{i\rightarrow j}_{t}$, can easily go into sub-optimal depth result, where we verify this in the experiment section in Figure 5.
+Therefore, apart from the spatial loss, [44] proposes to facilitate the Structure-from-Motion (SFM) to extract sparse depth information for direct depth supervision to provide a stronger supervision signal, but it is time-consuming and not straightforward.
 Different from [44], [27] enhances the depth estimation performance with spatio-temporal context that does not need the sparse depth from SFM but the performance is still limited.
@@
 We find that the nature of Gaussian splatting is scale-aware projection that could serve for the cross-view stereo constraint.
-We propose Gaussian splatting for projection in stage 1 for better scale-aware training.
+We propose Gaussian splatting for projection in stage 1 for better scale-aware training as follows.
 Gaussian Splatting for Projection (GSP): As illustrated in Figure 2, we adapt a depth network [52] to predict the Gaussian attributes in 2D grid space, where, apart from the original depth map, we also predict the scale map and rotation map.
 For each adjacent view, we first calculate the mask in the overlap region, then mask out one side of these overlap regions.
-Due to the presence of the other side’s overlap region, the unprojected 3D scene remains complete, which is critical for providing scale training, as indicated in the experiment section (Table 3).
+Due to the presence of the other side’s overlap region, the unprojected 3D scene remains complete if the depth map is predicted well.
+This mask-out step is critical for providing scale training, as indicated in the experiment section (Table 3).
 We then perform splatting rendering on the adjacent views to obtain the rendered image.
-If the depth map is accurately learned, the rendered image should resemble the original images, providing the necessary scale information for the entire pipeline.
-The process of acquiring the overlap mask is illustrated in Figure 3.
-The acquisition of the overlap mask is similar to volume rendering.
+If the depth map is accurately learned, the rendered image should resemble the original images, providing the necessary scale information for the joint training with 6D pose net.
+Overlap Mask: The process of acquiring the overlap mask is illustrated in Figure 3.
 We densely sample points along the ray in one view, and a pixel is considered part of the overlap region if more than one sampled 3D point falls within the adjacent view.
-The overlap mask is determined by the camera extrinsic and defined max depth (e.g., 80m in nuScenes).
+The overlap mask is only determined by the camera’s extrinsic and defined max depth (e.g., 80m in nuScenes).
 Note that in the DDAD dataset [13], we exclude regions with self-occlusion (such as parts of the vehicle body).
-Finally, we apply an erosion operation from OpenCV library [2] to the mask for purification.
+Besides, we apply the erosion operation from OpenCV [2] to the mask for purification.
 3.3 Fast rendering by Gaussian Splatting
-Inefficient performance in volume rendering: For 2D supervision (semantic and depth maps), previous methods [11, 54, 23, 37] employed volume rendering based on dense sampling.Although the final 3D voxel representation for modeling the 3D scene is much quicker than the original implicit representation [35], it remains time-consuming, particularly when incorporating semantic map rendering.
+Inefficient performance in volume rendering: For 2D supervision (semantic and depth maps), previous methods [11, 54, 22, 37] employed volume rendering based on dense sampling.
+Although the final 3D voxel representation for modeling the 3D scene is much quicker than the original implicit representation [35], it remains time-consuming, particularly when incorporating semantic map rendering.
 For example, in OccNeRF [54], the number of sampled points at a resolution of $180\times 320$ is 108,735,066.
-However, the target optimized points correspond to the vertices in the 3D voxel grid, which total $320\times 320\times 24=2,577,525$.
+However, the target optimized points correspond to the vertices in the 3D voxel grid, totaling $300\times 300\times 24=2,160,000$.
 This redundancy in densely sampled points helps optimization with volume rendering but is highly inefficient.
@@
 Then, we can optimize the attributes of the 3D Gaussian, such as semantic and opacity information.
-For example, in the empty space of the voxel, even though we have the vertices at that region during the splatting rendering, after the optimization, the network would predict the opacity as zero at these vertices then these vertices would not contribute any geometry or semantic information during the rendering.
-Since all the vertices are arranged in 3D voxel space with real-world 3D position $\mathcal{X}$, we use fixed scale $s$ and fixed rotation $r$ for each vertex for simplification.
+For example, in the empty space of voxel, even though we have the vertices at that region during the splatting rendering, after the optimization, the network would predict the opacity as zero at these vertices then these vertices would not contribute any geometry or semantic information during the rendering.
+Since all vertices are arranged in 3D voxel space with real-world 3D position $\mathcal{X}$, we can use fixed scale $s$ and fixed rotation $r$ for each vertex for simplification.
 This allows us to model the 3D scene by optimizing the rest Gaussian attributes (semantic and opacity).
@@
 $\mathcal{L}_{stage1}=\mathcal{L}_{temporal}+\mathcal{L}_{cross},$ (2)
-$\mathcal{L}_{stage2}=\mathcal{L}_{temporal}+\mathcal{L}_{semantic},$ (3)
-$(\mathcal{L}_{temporal},\mathcal{L}_{cross})=\frac{1-\operatorname{SSIM}\left(I_{t},\hat{I}_{t}\right)}{2}+(\beta)\left\|I_{t}-\hat{I}_{t}\right\|,$ (4)
+$\mathcal{L}_{stage2}=\mathcal{L}_{temporal}+\lambda\mathcal{L}_{semantic},$ (3)
+$(\mathcal{L}_{temporal},\mathcal{L}_{cross})=\frac{1-\operatorname{SSIM}\left(I_{t},\hat{I}_{t}\right)}{2}+\beta\left\|I_{t}-\hat{I}_{t}\right\|,$ (4)
 where $I_{t}$ and $\hat{I}_{t}$ refer to the target image and the corresponding synthesized image, respectively.
@@
 In contrast, $\hat{I}_{t}$ in the cross-view photometric loss $\mathcal{L}_{cross}$ is derived from our proposed cross-view Gaussian splatting method.
-$\mathcal{L}_{semantic}$ is 2D semantic loss.
-The $\beta$ is set to 0.15 to balance the two loss terms.
-For $\mathcal{L}_{temporal}$ and $\mathcal{L}_{semantic}$, we recommend [54] for further details.
+$\mathcal{L}_{semantic}$ is 2D semantic loss with balanced weight $\lambda=0.02$.
+For $\mathcal{L}_{cross}$ and $\mathcal{L}_{temporal}$, we set $\beta$ is set to 0.15 for weight balance the same as [54].
 Figure 3: Overlap mask in nuScenes [3] and DDAD [13].
+4 Experiment
+4.1 Tasks, datasets, and metric
+nuScenes [3]: For 3D occupancy estimation, we utilize annotations from Occ3D [41].
+For fair comparison, we use the 2D pseudo semantic map provided by OccNeRF [54] for training.
+We measure 3D occupancy estimation performance using the mean Intersection over Union (mIoU) metric.
+For depth estimation, we set the perception range in [-80m, -80m, -1m, 80m, 80m, 6m], while we clamp the ground truth to a range of 0.1m to 80m for evaluation, consistent with OccNeRF and SurroundDepth [44].
+We evaluate depth maps using error metrics (Abs Rel, Sq Rel, RMSE, RMSE log) and threshold accuracy metrics ($\delta$).
+DDAD [13]: Though we do not have the 3D occupancy labels on DDAD dataset, we can present qualitative results thanks to our fully self-supervised 3D occupancy estimation setting.
+We obtain the 2D pseudo semantic labels for training following OccNeRF pipeline in nuScenes dataset.
+For depth estimation, we clamp the depth range within 0.1m and 200m for evaluation, consistent with SurroundDepth [44].
 Method GT Occ. GT Pose mIoU* mIoU
@@
 OccFormer [57] $\times$ 22.39 21.93
-RenderOcc [37] $\times$ 24.53 23.93
 TPVFormer [21] $\times$ 28.69 27.83
 CTF-Occ [41] $\times$ 29.54 28.53
+RenderOcc [37] $\times$ $\times$ 24.53 23.93
 SimpleOcc [11] $\times$ 7.99 7.05
@@
 GaussianOcc $\times$ $\times$ 11.26 9.94
-Table 1: 3D occupancy prediction performance on the Occ3D-nuScenes dataset in mIoU metric.
+Table 1: 3D occupancy comparison on the Occ3D dataset with mIoU metric.
 Since ‘other’ and ‘other flat’ classes are the invalid prompts for open-vocabulary models, we also calculate ‘mIoU*’ as the result ignoring the classes that do not consider these two classes during evaluation, while ‘mIoU’ is the original result.
 GT Occ.
-means using the ground truth occupancy label for the supervision.
-GT Pose is the ground truth pose from the sensor for self-supervised geometry learning.
-The full table is placed in the supplementary material with each category result.
+means using the ground truth occupancy label for supervision.
+GT Pose is the ground truth 6D ego pose from the sensor for self-supervised geometry learning.
 Method GT pose Occ. Abs Rel Sq Rel RMSE RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
@@
 Figure 4: Visualization of the render depth map and 3D occupancy prediction on the nuScenes and DDAD datasets.
-4 Experiment
-4.1 Tasks, datasets, and metric
-nuScenes [3]: For 3D occupancy estimation, we utilize annotations from Occ3D-nuScenes [41], which includes 600 scenes for training and 150 scenes for evaluation.
-Each sample has a perception range of [-40m, -40m, -1m, 40m, 40m, 5.4m] with a voxel size of 0.4m.
-For fair comparison, we use the 2D semantic map provided by OccNeRF [54].
-We measure 3D occupancy estimation performance using the mean Intersection over Union (mIoU) metric.
-For depth estimation, we set the perception range in [-80m, -80m, -1m, 80m, 80m, 6m], while we clamp the ground truth to a range of 0.1m to 80m for evaluation, consistent with OccNeRF and SurroundDepth [44].
-We evaluate depth maps using error metrics (Abs Rel, Sq Rel, RMSE, RMSE log) and threshold accuracy metrics ($\delta$).
-DDAD [13]: Due to the lack of ground truth labels, we only present qualitative results for the fully self-supervised 3D occupancy estimation setting.
-Following the procedure on the nuScenes dataset, we obtain the 2D semantic labels according to the OccNeRF pipeline.
-For depth estimation, we also set the perception range in [-80m, -80m, -1m, 80m, 80m, 6m], but clamp the ground truth to a range of 0.1m to 200m for evaluation, consistent with SurroundDepth [44].
 Scale-aware training in [44, 27] Scale-aware training by ours RMSE
@@
 Mask represents using the mask-out strategy before the unprojection.
-Erode means the erode process to the binary overlap mask and Refine is the refinement of depth estimation network with 2 epochs by fixing the 6D pose network.
+Erode means the erode process to the binary overlap mask and Refine is the refinement of depth estimation network with 2 epochs by fixing the 6D pose net.
 4.2 Implementation details
 Network details: For U-Net architecture, we adapt New-CRFs [52] to predict the Gaussian attributes, which is based on the Swin Transformer [33].
-The 6D pose network is the same as that used in SurroundDepth [44].
+The 6D pose net is the same as that used in SurroundDepth [44].
 For the 2D-to-3D lifting, we follow the approach used in SimpleOcc [11].
@@
 For occupancy estimation, we use the same network as OccNeRF [54] to ensure a fair comparison.
-Training details: We propose a two-stage training process for fully self-supervised 3D occupancy estimation.
-In stage 1, we jointly train the depth estimation network and the 6D pose network, where we train the models for 8 epochs on the nuScenes dataset and 12 epochs on the DDAD dataset.
+Training details: We propose a two-stage training for fully self-supervised 3D occupancy estimation as indicated in Figure 2.
+In stage 1, we jointly train the depth estimation network and the 6D pose net, where we train the models for 8 epochs on the nuScenes and 12 epochs on the DDAD.
 In stage 2, we train the 3D occupancy network in a self-supervised manner with the 6D pose predicted from stage 1 rather than the ground truth pose used in OccNeRF [54].
-We train the models for 12 epochs on both the nuScenes and DDAD datasets.
+We train the models for 12 epochs on both the nuScenes and DDAD.
 The optimizer and learning rate adjustment strategy follow those used in SimpleOcc [11] and OccNeRF [54].
-More implementation details (including code) are provided in the supplementary material.
 4.3 Main results
-3D occupancy estimation in nuScenes: As shown in Table 1, the proposed GaussianOcc achieves the best performance compared to other self-supervised methods.
-In particular, we do not require ground truth occupancy labels and ground truth 6D poses for training, thanks to the predicted 6D pose from the stage 1.
-In addition, we also achieve the good result on the RayIoU metric [31], with further details provided in the supplementary material.
-3D occupancy estimation in DDAD: Thanks to our fully self-supervised learning setting, we are able to perform 3D occupancy estimation on the DDAD dataset [13].
-To the best of our knowledge, our work is the first to introduce 3D occupancy estimation on this dataset, although we cannot provide quantitative results due to the lack of 3D occupancy label.
+3D occupancy estimation in nuScenes: In Table 1, the proposed GaussianOcc achieves the best performance compared to other self-supervised methods.
+In particular, we do not require ground truth occupancy labels and ground truth 6D ego pose for training, thanks to the predicted 6D pose from the stage 1.
+Note that RenderOcc [37] does not require the 3D occupancy label, but it is not a self-supervised method since it uses the ground truth depth map and semantic map for the 2D supervision; it could be regarded as a weakly-supervised method.
+In addition, we also achieve the best result on the RayIoU metric [31], with further details provided in the supplementary material.
+3D occupancy estimation in DDAD: To the best of our knowledge, our work is the first to achieve 3D occupancy estimation on this dataset, thanks to our fully self-supervised learning setting.
 We present visualization results in Figure 4.
-More results can be found in the supplementary material.
-Depth estimation: We present a comparison of depth estimation results in Table 2 for both the nuScenes and DDAD datasets.
-In stage 1, GaussianOcc $\ddagger$ achieves top performance on the nuScenes dataset and delivers competitive results on the DDAD dataset.
+As highlighted by the red rectangle, the sky region has a short-range depth value, but this does not appear in the rendered 3D occupancy estimation map thanks to the parameterized coordinate design of OccNeRF [54].
+Depth estimation: We present a comparison of depth estimation results in Table 2 for both the nuScenes and DDAD.
+In stage 1, GaussianOcc $\ddagger$ achieves top performance on the nuScenes dataset and delivers competitive results on the DDAD.
 It is important to note that methods such as SurroundDepth [44] and SA-FSM [50] rely on third-party module for the sparse depth supervision.
@@
 In stage 2, which involves depth estimation from rendering, our method also achieves competitive results compared to those trained with ground truth poses.
-Besides, we observe that the rendered depth in stage two outperforms the depth results from stage 1 on the nuScenes dataset, whereas the opposite is true for the DDAD dataset.
+Besides, we observe that the rendered depth in stage two outperforms the depth results from stage 1 on the nuScenes, whereas the opposite is true for the DDAD.
 This discrepancy might be attributed to differences in perception range—80 meters in nuScenes versus 200 meters in DDAD.
 Moreover, an interesting phenomenon is that the semantic information is helpful for the depth estimation as indicated in GaussianOcc $\dagger$ whereas it worsens the result in OccNeRF $\dagger$.
-This shows the difference between the Gaussian splatting rendering and volume rendering.
+This phenomenon can be attributed to the biased sampling strategy of OccNeRF, where only 25% of the sample points are used for faster semantic map rendering compared to depth map rendering.
+In contrast, our proposed Gaussian splatting method, which renders directly from the voxel vertices, eliminates this issue.
+Figure 5: The comparison of the depth map and its synthesis overlap image with (1) direct bilinear interpolation cross-view synthesis [44] and (2) our cross-view Gaussian splatting synthesis.
 4.4 Ablation Study
-Scale-aware training: To demonstrate the effectiveness of the proposed scale-aware training using Gaussian Splatting, we conducted experiments comparing our method with existing approaches [44, 27] on Table 3.
+Scale-aware training in Stage 1: We compare our method with existing approaches [44, 27] on Table 3 to demonstrate the effectiveness of the proposed scale-aware training using Gaussian Splatting.
 For a more fair comparison, we also implemented scale-aware training on [44, 27] that uses the same depth estimation network as ours [52], noted as $\checkmark$.
@@
 Our proposed scale-aware training method is better than [44, 27].
-Specifically, we observed that a naive implementation of Gaussian splatting without a masking strategy for cross-view rendering is ineffective because it would lead to a sub-optimal solution that the rendered image is still from the current view.
+Specifically, we observed that a naive implementation of Gaussian splatting without a mask-out strategy for cross-view rendering is ineffective because it would lead to a sub-optimal solution that the rendered image is still from the current view.
 Besides, to enhance performance, we introduced an erosion operation on the binary mask to purify it, excluding regions that may fall outside the overlap area.
 This step ensures better alignment during training.
-Finally, we refined the depth estimation by fixing the 6D pose network and disabling the cross-view loss.
+Finally, we refined the depth estimation by fixing the 6D pose net and disabling the cross-view loss.
 This refinement helps reduce artifacts at the edges of the overlap region.
-We present the cross-view splatting render result in Figure 5, where the rendering result in overlap regions is reasonable in the correct depth estimation from cross view that verifies our design.
-6D pose learning and training strategy: Considering the GT pose from the sensor is imperfect as a lack of vertical movement [36].
-We further evaluate the learned 6D pose quality by the self-supervised depth result at stage 2, where the experiments on Table 4 use the same network but with different pose constraints.
-Our result is competitive compared with GT pose, indicating that the predicted pose is of high quality.
-Furthermore, one stage training was ineffective in using the 6D pose from the jointly trained with 6D pose network because the cross-view loss in the 3D voxel space led to local optimization, which failed to generalize predictions to non-overlapping regions., as shown in Figure 6.
-However, this issue did not arise in the depth maps generated by the 2D decoder in stage 1, which is why we need the two-stage training process.
-Besides, our result is also better than [27], which is consistent the depth result on Table 3.
-Volume rendering and Splatting rendering: In Table 5, we compare the performance of volume rendering (VR) and splatting rendering (SR) in the voxel space.
-For each 3D Gaussian, we set a uniform scale $s$ for all Gaussians, considering the well-arranged positions of vertices within the voxel grid.
-Since the scale $s$ is consistent across all three dimensions, we set the rotation $\mathbf{R}$ as the identity matrix.
-We first explored the optimal scale and found that a scale of 0.1 produced the best results.
-Notably, the performance of VR and SR was very close when using this scale.
-The render depth maps are shown in Figure 6.
-The depth map by the volume rendering is smoother thanks to dense sampling and the depth maps from splatting tendering have the graininess effect, especially at the small scale factor.
-Pose and render types ablation study in occupancy task: In Table 6, we present an ablation study to evaluate the impact of using ground truth versus learned poses in the occupancy task, with training conducted using different rendering methods.
-The results show that splatting rendering achieves superior performance in both occupancy metrics (mIoU*) and depth metrics.
-One observation is that depth estimation results with the semantic learning in volume rendering are significantly worse than those in Table 2 whereas the proposed splatting rendering method maintains consistent performance, which suggests the splatting rendering in voxel grid contributes better 3D geometry.
-Rendering efficiency analysis: We analyze rendering efficiency in Table 7.
-The results show that volume rendering consumes 5 times more rendering time compared to splatting rendering.
-As resolution increases, both volume rendering time and GPU consumption rise significantly.
-In contrast, splatting rendering shows no significant increase in computational cost with higher resolutions, highlighting its efficiency and scalability.
-Bonus of the fully self-supervised setting: The fully self-supervised setting of our method could be a general pretraining solution for supervised learning.
-After the self-supervised training on the DDAD and nuScenes datasets, we further finetune the model with the 3D occupancy label from Occ3D-nuScenes [41].
-As shown in Table 8, experiments with self-supervised pretraining outperform the baseline.
-In particular, we find that pretraining on nuScenes is better than the DDAD dataset, which may own to the domain gap factors, such as differences in the scenarios (RGB images) and sensor configurations (camera extrinsics).
+Visualization: In Figure 5, we can see that the depth map is easily be suboptimal depth estimation in near overlap region with the training of (1) direct cross-view interpolation constrain in [44].
+Specifically, though the region with yellow rectangle has the incorrect large depth value, the synthesis region still presents a reasonable result that would lead to sub-optimal training.
+Thanks to our cross-view Gaussian splatting design analysis above, our method shown in (2) does not have the sub-optimal situation.
+Figure 6: The comparison for the depth map in the different setting, corresponding to the training strategy in Table 4 and rendering type in Table 5.
 Pose type Abs Rel Sq Rel RMSE $\delta<1.25$
@@
 SR ($s$ = 0.15) 0.217 3.406 7.204 0.763
+SR ($s$ = learnable) 0.212 3.248 7.112 0.771
 Table 5: Comparison of the render result between the volume rendering (VR) [54] and splatting rendering (SR, Ours) on depth estimation task [3].
-The coefficient $s$ means the scale of the 3D Gaussian.
+The coefficient $s$ means the scale of the 3D Gaussian and learnable means the scale is learnable by the Sigmoid function.
 We use GT pose for this set of the ablation study.
@@
 The render time is calculated from surround 6 images.
-”N/A” indicates out-of-memory errors running in NVIDIA A 100 (40 GB).
+"N/A" indicates out-of-memory errors running in NVIDIA A 100 (40 GB).
 Training time is averaged per epoch.
-Method Pretraining setting mIoU
-Baseline None 37.29
-Self-supervised pretraining DDAD 37.40
-nuScenes 38.45
-Table 8: The study on SimpleOcc [11] with fully self-supervised pretraining.
-The baseline is directly training the model with 3D occupancy labels.
-The self-supervised pretraining is conducted on DDAD and nuScenes and then finetuned the model with 3D occupancy label.
-The number with bold typeface means the best.
-Figure 5: Visualization of the depth map, splatting render image and depth map from cross view in stage 1.
-Figure 6: The comparison for the depth map in the different setting, corresponding to the training strategy in Table 4 and rendering type in Table 5.
+6D pose learning and training strategy: Considering the 6D ego pose from the sensor is imperfect as a lack of vertical movement [36].
+We further evaluate the learned 6D pose quality by self-supervised depth result at stage 2, where the experiments on Table 4 use the same network but with different pose constraints.
+Our result is competitive compared with GT pose, indicating that the predicted pose is of high quality.
+Furthermore, one stage training was ineffective in using the 6D pose from the jointly trained with 6D pose net because the cross-view loss in the 3D voxel space led to local optimization, which failed to generalize predictions to non-overlapping regions, as shown in Figure 6.
+However, this issue did not occur in the depth maps produced by the 2D decoder in Stage 1, highlighting the necessity of the two-stage training.
+Besides, our result is also better than [27], which is consistent the depth result on Table 3.
+Volume rendering and Splatting rendering: In Table 5, we compare the performance of volume rendering (VR) and splatting rendering (SR) in the voxel space.
+For each 3D Gaussian, we first set a uniform scale $s$ for all Gaussians, considering the well-arranged positions of vertices within the voxel grid.
+Since the scale $s$ is consistent across all three dimensions, we set the rotation $\mathbf{R}$ as the identity matrix.
+We found that a scale of 0.1 produced the best results, closely matching the VR at this scale.
+Besides, by applying a learnable scale through the output of a Sigmoid function and clamping the maximum scale to 0.12, we achieved the highest performance.
+The render depth maps are shown in Figure 6.
+The depth map by the volume rendering is smoother thanks to dense sampling and the depth maps from splatting rendering have the graininess effect, especially at the small scale factor (0.05).
+Pose and render types ablation study in occupancy task: In Table 6, we present an ablation study to evaluate the impact of using ground truth versus learned poses in the occupancy task, with training conducted using different rendering methods.
+The results show that splatting rendering achieves superior performance in both occupancy metrics (mIoU*) and depth metrics.
+One observation is that depth estimation results with the semantic learning in volume rendering are significantly worse than those in Table 2 whereas our proposed splatting rendering method maintains consistent performance, which suggests the splatting rendering in voxel grid contributes better 3D geometry.
+Rendering efficiency analysis: We analyze rendering efficiency in Table 7.
+The results show that volume rendering consumes 5 times more rendering time compared to splatting rendering.
+As resolution increases, both volume rendering time and GPU consumption rise significantly.
+In contrast, splatting rendering shows no significant increase in computational cost with higher resolutions, highlighting its efficiency and scalability.
 5 Conclusion
@@
 Additionally, the proposed Gaussian splatting in voxel grids outperforms volume rendering in 3D occupancy estimation while reducing computational cost.
-As a bonus, the fully self-supervised setting of the proposed method could serve as the general pretrain solution to enhance the 3D occupancy estimation performance.
-We validate our method on the nuScenes and DDAD datasets, demonstrating its strong generalization ability to surround view scenes.
-6 More implementation details
+Acknowledgments This work was supported in part by the JSPS, KAKENHI under Grant Number 22H03609, JST, FOREST under Grant Number JPMJFR206S.
+Wanshui Gan was also supported by RIKEN JRA Program.
+Big thanks for Xiaoyu Dong for helping proofreading the manuscript.
+Appendix
+Appendix A More implementation details
 The detailed parameter setting in Gaussian attributes estimation network [52].
@@
 We do not predict the color defined by SH coefficients $c$, while we directly use the source RGB image as the color map the same as in [58].
-The detailed parameter setting in voxel grid splatting rendering.
-We chose a fixed scale for each grid vertex to ensure a well-arranged structure that accurately models the 3D space.
+The detailed parameter setting in voxel grid splatting rendering for semantic rendering.
+For semantic rendering, we chose a fixed scale for each grid vertex to ensure a well-arranged structure that accurately models the 3D space.
 If we use a learnable scale, it may lead to a situation where the scale is small but the opacity is large, which may not be captured in the rendered depth map and semantic map but could still affect the 3D occupancy result.
-Therefore, using a fixed scale is simple and sufficient for optimization, as demonstrated in the results presented in the main paper.
+Therefore, using a fixed scale is simple and sufficient for optimization, as demonstrated in the results presented in the main paper (Table 5) that the performance is close to the learnable scale.
 Since the scale $s\in\mathbb{R}^{3}_{+}$ are identical for both three dimensions, we do not need to predict the rotation $r\in\mathbb{R}^{4}$ and set it with the identical matrix is sufficient.
@@
 where $M$ is the valid pixel, $\hat{d}$ is the ground truth depth and $d^{*}$ is the predicted depth.
-7 More experiment results
+Figure 7: One-stage training analysis. This is the visualization of the overlap mask and the rendered depth map with one-stage training.
+Appendix B More experiment results and analysis
+Why we need two-stage training.
+We made extensive efforts to develop one-stage training that directly applies the cross-view loss to the rendered depth map but were unsuccessful.
+As suggested in Figure 7, the cross-view supervision signals are effective only in overlapping regions.
+The rendered depth map learned from 3D CNN has lower generalization ability in non-overlap regions compared with the decoder depth learned by the 2D CNN, which led to local minima.
 mIoU metric. Due to the limited space in the main paper, the full table of mIoU results is presented in Table 9 for reference.
@@
 Since GaussianOcc and OccNeRF utilize the same network architecture, they share the same inference time when the rendering process is not taken into account.
-More visualization. We provide more visualization for nuScenes dataset in Figure 7 and DDAD dataset in Figure 8. Please check the attached videos for sequence visualization.
-More analysis on different supervision types.
-In Figure 9, we present visualizations of different supervision types.
+More visualization. We provide more visualization for nuScenes dataset in Figure 8. Please check the videos for sequence visualization in https://github.com/GANWANSHUI/GaussianOcc.git.
+More analysis on 3D occupancy and depth map result on different supervision types.
+3D occupancy analysis: In Figure 9, we present visualizations of different supervision types.
 These visualizations highlight key differences in the results for the invisible regions (marked with red rectangles) and the rendered depth quality (marked with green rectangles).
@@
 Conversely, Experiment (3) demonstrates a consistent tendency to classify invisible regions as man-made structures, likely because the surrounding environment predominantly consists of man-made elements.
-Additionally, the depth maps in the green-highlighted areas show that our Gaussian splatting rendering produces higher-quality depth predictions compared to volume rendering, consistent with the quantitative results in Table 6 of the main paper.
-Since Experiments (1) and (2) do not involve rendering-based training, they fail to produce reasonable depth predictions.
-8 Limitation and future work
+Render depth map analysis: In Tables 2 and 6 of main paper, we observe an interesting phenomenon that the semantic information is helpful for the depth estimation with our GaussianOcc whereas it worsens the result in OccNeRF.
+In Figure 9, we visualize the depth map and highlight with green rectangles that our Gaussian splatting rendering produces higher-quality depth predictions compared to volume rendering.
+This should be concluded to the biased sampling strategy of OccNeRF, where only 25% of the sample points are used for faster semantic map rendering compared to depth map rendering.
+Here is the piece of the code in OccNeRF [54].
+In contrast, our proposed Gaussian splatting method, which renders directly from the voxel vertices, eliminates this issue.
+At last, since Experiments (1) and (2) do not involve rendering-based training, they fail to produce reasonable depth predictions.
+Render resolution Render time (s) with different voxel resolutions (Gaussians number)
+$180\times 320$ $16\times 200\times 200$ $24\times 300\times 300$ $32\times 512\times 512$
+VR $\approx$ 0.50 $\approx$ 0.85 $\approx$ 1.52
+SR $\approx$ 0.06 $\approx$ 0.17 $\approx$ 0.44
+Table 8: Comparison of rendering efficiency under different Gaussians number between volume rendering (VR) [54] and splatting rendering (SR, Ours).
+Gaussians number and its related render time: (1) In stage 1, Gaussians number depends on the depth map resolution from the 2D decoder, where each pixel is a Gaussian primitive after unprojection.
+We use the depth map resolution in $224\times 352$, resulting in 78,848 Gaussian primitives in one image.
+In stage 2, Gaussians number depends on the voxel resolution, where each voxel grid is a Gaussian primitive.
+In Table 7 of the main paper, we follow the voxel resolution the same as OccNeRF [52] in $24\times 300\times 300$, resulting in 2,160,000 Gaussian primitives.
+(2) We revealed the rendering time under different render image resolutions compared with volume rendering in Table 7 of the main paper.
+We conducted the extra experiment for render time comparison under the same render image resolution ($180\times 320$) but with different Gaussians number (voxel resolutions) as shown in Table 8.
+From Table 7 of the main paper and Table 8, we observe: (1) The render time of splatting rendering (SR) is mainly affected by the Gaussians number, not the render image resolution.
+(2) SR is 3–8 times faster than volume rendering (VR) across different voxel settings.
+Bonus of the fully self-supervised setting: The fully self-supervised setting of our method could be a general pretraining solution for supervised learning.
+After the self-supervised training on the DDAD and nuScenes datasets, we further finetune the model with the 3D occupancy label from Occ3D [41].
+As shown in Table 11, experiments with self-supervised pretraining outperform the baseline.
+In particular, we find that pretraining on nuScenes is better than the DDAD dataset, which may own to the domain gap factors, such as differences in the scenarios (RGB images) and sensor configurations (camera extrinsics).
+Appendix C Limitation and future work
 The proposed method achieves reasonable predictions in most scenes; however, we observe that some cases still present challenges, as shown in Figure 10.
@@
 We believe that this problem could be mitigated with better 2D semantic maps for supervision, which warrants further investigation.
-The proposed method is for the surround view setting which is not suitable for the monocular images and the two-stage training procedure is a little bit time-consuming.
-Therefore, in the future, we will further investigate to apply the proposed method for one-stage training.
-Additionally, in stage 1, we leverage the spatial cross-view constraint through the proposed Gaussian splatting method.
+The proposed method is for the surround view setting which is not suitable for the monocular images.
+Additionally, in stage 1, we leverage the spatial cross-view constraint for scale-aware training through the proposed Gaussian splatting method.
 In the future, we aim to explore its potential benefits for temporal view synthesis as well.
@@
 Figure 8: The visualization of the render depth map and 3D occupancy prediction on nuScenes dataset.
-Figure 8: The visualization of the render depth map and 3D occupancy prediction on DDAD dataset
 Figure 9: The visualization of the different supervision types (1-4) comparison on nuScenes dataset.